### v1

# Semi-Blind Strategies for MMSE Channel Estimation Utilizing Generative Priors

Abstract
This paper investigates semi-blind channel estimation for massive multiple-input multiple-output (MIMO) systems. To this end, we first estimate a subspace based on all received symbols (pilot and payload) to provide additional information for subsequent channel estimation. We show how this additional information enhances minimum mean square error (MMSE) channel estimation. Two variants of the linear MMSE (LMMSE) estimator are formulated, where the first one solves the estimation within the subspace, and the second one uses a subspace projection as a preprocessing step. Theoretical derivations show the superior estimation performance of the latter method in terms of mean square error for uncorrelated Rayleigh fading. Subsequently, we introduce parameterizations of this semi-blind LMMSE estimator based on two different conditional Gaussian latent models, i.e., the Gaussian mixture model and the variational autoencoder. Both models learn the underlying channel distribution of the propagation environment based on training data and serve as generative priors for semi-blind channel estimation. Extensive simulations for real-world measurement data and spatial channel models show the superior performance of the proposed methods compared to state-of-the-art semi-blind channel estimators with respect to the MSE.
Index Terms:
Semi-blind channel estimation, Gaussian mixture model, variational autoencoder, measurement data.
I Introduction
Accurate channel state information (CSI) is crucial for achieving the expected high data rates promised by multiple-input-multiple-output (MIMO) systems [2, 3, 4]. The CSI describes the communication link between transmitter and receiver, characterized by its time-varying and frequency-selective nature, which is prone to rapid changes making the task of channel estimation complex [5]. As accurate channel estimates are essential for the successful transmission of data, it is placed at the center of several research efforts [6, 7].
The most widely adopted methods in wireless communication utilize known training or pilot symbols transmitted across the channel using some of the radio resource blocks [8]. Afterward, the receiver uses the observed signals to determine a reliable CSI estimate. As the number of pilots scales with the number of users, the spectral efficiency decreases for higher number of users as less symbols are available for transmitting data. To enhance channel estimation without increasing the number of pilot symbols, various methods have been developed that leverage the information embedded in the observed data symbols at the receiver to infer channel characteristics [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These methods exploit structure and redundancy within the transmitted data and yield more accurate CSI estimates.
The benefit of semi-blind channel estimation was first studied in [9], where Cramer-Rao bounds (CRBs) for blind, semi-blind, and training-based channel estimation were investigated in the context of single-input-multiple-output (SIMO) systems. In [10, 11], semi-blind channel estimation schemes based on maximum likelihood (ML) estimation were introduced. The asymptotic performance of the respective estimators was also studied in [10] for infinitely long data sequences. Another asymptotic behavior, where the number of antennas grows to infinity, was studied in [12]. Here, the authors identified two interference components in semi-blind channel estimation, which do not vanish even for large numbers of antennas. Early work on improving the least squares (LS) estimator using a semi-blind algorithm was conducted in [13]. The authors in [14] propose to use partially decoded data for channel estimation. Furthermore, in [15], semi-blind and blind channel estimation was studied to enhance the maximum a-posteriori (MAP) channel estimates in massive MIMO systems. These findings are based on favorable propagation, which only holds for large antenna arrays deployed at the base station (BS). In [16], two semi-blind channel estimators based on the expectation maximization (EM) algorithm are studied. The assumption of a Gaussian distribution for the data was verified, leading to a closed-form solution for the E-step. Another iterative framework optimizing the likelihood based on message passing (MP) is used in [17, 18] and references therein. In [19], a data-aided iterative scheme is proposed for orthogonal time frequency space (OTFS) systems by employing affine-precoded superimposed pilots. The performance improvement achieved with these iterative approaches generally requires computational costly updates. A low complexity iterative LS channel estimation algorithm is proposed in [20] for a massive MIMO turbo-receiver. In [21], the concept of semi-blind channel estimation was adapted for time domain synchronous-orthogonal frequency division multiplexing (OFDM) systems, where in addition to the pseudo noise sequence in the guard interval, the sent OFDM data symbols are exploited for the channel estimation. In [22], a framework was introduced for choosing reliable decoded data symbols, which can be interpreted as additional pilots. Similarly, reliably detected symbols are used for a semi-data-aided channel estimation in [23]. In [24], peak-power carriers in an OFDM system are selected to eliminate the need to determine reliable data symbols at the receiver. Recently, a diffusion model-based approach for joint channel estimation and detection was proposed in [25], where a diffusion process is constructed that models the joint distribution of the channels and symbols given noisy observations.
In this article, we focus on pilot-based estimators which minimize the MSE and investigate how these estimators can be extended for the semi-blind case. Notably, in [26] a subspace formulation of the MMSE estimator was used to mitigate pilot contamination in massive MIMO systems. The MMSE estimator is known to be the conditional mean estimator (CME) [27, Ch. 10], which, in general, is intractable and can not be computed in closed form. Recently, powerful approximations based on machine learning were presented in [28, 29, 30, 31, 32]. The benefit of machine learning is to enhance the task at hand by using prior information captured during the learning stage. For a given BS cell environment, the probability density function (PDF) representing potential user channels can be considered valuable prior information. Since this true underlying distribution is unknown, machine learning methods rely on a representative data set, which is assumed to be available at the BS. Based on this data set the PDF of the user channels can be learned. The first proposal of using a Gaussian mixture model (GMM) to formulate an estimator was done in [28] for the case of image processing. The approaches in [30, 31, 32] build on that by constructing a conditionally Gaussian latent model (CGLM) for the PDF of a BS cell environment. The learned CGLM not only enables MMSE channel estimation in [30, 31, 32] but can also be used for e.g., a limited feedback scheme as in [33]. In this work, we propose to utilize CGLMs to parameterize the CME in the semi-blind setting.
The contributions of this work are summarized as follows:
  • We introduce two variants of the linear MMSE (LMMSE) estimator incorporating subspace knowledge provided by the payload data symbols. First, we depict how the LMMSE channel estimator can solve a subspace estimation problem [26]. As an alternative, we propose a projection method that is computationally more efficient since it allows for the pre-calculation of LMMSE filters.
  • With theoretical derivations we show the superior MSE performance of the proposed projection method in the case of uncorrelated Rayleigh fading and perfect subspace knowledge.
  • We show how the GMM [30] and variational autoencoder (VAE) [32], instances of the class of CGLMs, can be used to parameterize the semi-blind LMMSE estimator.
  • Extensive simulations on different datasets, consisting of typical massive MIMO systems with multiple users and including real-world measurement data, show the superior performance of our proposed methods compared to state-of-the-art semi-blind channel estimation algorithms with respect to the MSE.
Preliminary results were presented in [1] and extended to the multi-user MIMO case in [34], which we extend further in the following aspects. The theoretical analyses in Section III enhance the foundation of the proposed semi-blind channel estimation strategies and provide analytic insights into the superior performance of the proposed projection method. We extend our concept of semi-blind MMSE channel estimation to the whole class of CGLMs, providing a more general framework to parameterize the semi-blind LMMSE estimator. Finally, we provide more comprehensive simulation results to show the strengths of our proposed strategies.
[…]
II System and Channel Model
We consider a multi-user uplink system with $J$ single-antenna users and a BS equipped with $M$ receive antennas. The received signal vector at time instance $n$ is then
${\bm{y}}(n)$ $={\bm{H}}{\bm{x}}(n)+{\bm{n}}(n),\quad n=1,...,N,$ (1)
where ${\bm{x}}(n)=[x_{1}(n),...,x_{J}(n)]^{\mathrm{T}}\in\mathbb{C}^{J}$ and ${\bm{n}}(n)\in\mathbb{C}^{M}$ denote the signal sent by each of the $J$ users and the additive noise, respectively, whereas ${\bm{H}}=[{\bm{h}}_{1},...,{\bm{h}}_{J}]$ contains the individual channels of the users ${\bm{h}}_{j}\in\mathbb{C}^{M}$. The case of multiple antennas at the users can be transformed into (1) by viewing each active stream as a different user. The corresponding channel ${\bm{h}}_{j}$ can then be seen as the effective channel. For further details on semi-blind channel estimation in multi-user MIMO, we refer the reader to [34]. For the task of channel estimation, we consider a channel coherence interval larger than the number of snapshots $N$, i.e., the channels are constant over all snapshots. We assume that the noise is Gaussian with ${\bm{n}}(n)\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M})$.
In conventional channel estimation schemes, each user’s signals include $N_{p}$ uplink pilots. These pilots are known to the BS. Hence, the received observations at the BS side are
${\bm{Y}}=\left[{\bm{Y}}_{p}^{\prime},{\bm{Y}}_{d}\right]={\bm{H}}\left[{\bm{P}},{\bm{D}}\right]+{\bm{N}}={\bm{H}}{\bm{X}}+{\bm{N}},$ (2)
where ${\bm{Y}}\in\mathbb{C}^{M\times N}$, ${\bm{Y}}^{\prime}_{p}\in\mathbb{C}^{M\times N_{p}}$, ${\bm{Y}}_{d}\in\mathbb{C}^{M\times N-N_{p}}$, ${\bm{P}}\in\mathbb{C}^{J\times N_{p}}$, and ${\bm{D}}\in\mathbb{C}^{J\times N-N_{p}}$ denote all received observations, received pilot observations, received payload data observations, sent pilots, and sent payload data symbols, respectively. In order to fully illuminate the channels, the number of pilots is, at minimum, the number of users $N_{p}\geq J$, and orthogonal pilots are used. We set $N_{p}=J$, and utilize discrete Fourier transform (DFT) pilot sequences. After decorrelating the orthogonal pilot sequences the received pilot observations simplify to
${\bm{Y}}_{p}={\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}={\bm{H}}{\bm{P}}{\bm{P}}^{\mathrm{H}}+{\bm{N}}{\bm{P}}^{\mathrm{H}}={\bm{H}}+{\bm{N}}_{p},$ (3)
where ${\bm{N}}_{p}$ has the same statistics as ${\bm{N}}$ and, hence, we can omit the subscript. This enables to consider channel estimation from a per user perspective in the subsequent discussions. For reasons of simpler readability, the index for the respective user is, therefore, no longer given in the following. Consequently, we denote the pilot observation of a user as
${\bm{y}}_{p}={\bm{h}}+{\bm{n}},$ (4)
with ${\bm{n}}\sim\mathcal{N}_{\mathbb{C}}(0,{\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M})$.
II-A Spatial Channel Model
We consider a spatial channel model based on [35], where the channel vectors are considered as conditionally Gaussian distributed [29]
${\bm{h}}\mid{\bm{\delta}}\sim\mathcal{N}_{\mathbb{C}}\left(\bm{0},{\bm{C}}_{\bm{\delta}}\right),$ (5)
based on a set of parameters ${\bm{\delta}}$, which describe the directions and properties of the multi-path propagation clusters. The main angles are drawn independently from a uniform distribution in $[0,2\pi]$ and the path gains are independent zero-mean Gaussians. The spatial covariance matrix is given by
${\bm{C}}_{\bm{\delta}}=\int_{-\pi}^{\pi}g(\vartheta,{\bm{\delta}}){\bm{a}}(\vartheta){\bm{a}}^{\mathrm{H}}(\vartheta)\mathrm{d}\vartheta,$ (6)
where $g(\vartheta,{\bm{\delta}})$ is the power density consisting of a weighted sum of Laplace densities, which have standard deviations $\sigma_{\text{AS}}$ corresponding to the angular spread of the propagation clusters. The BS employs a uniform linear array (ULA) with $M=64$ antennas and $\lambda/2$ spacing. The steering vector is then given as
${\bm{a}}(\vartheta)=\frac{1}{\sqrt{M}}\left[1,\mathrm{e}^{-\mathrm{j}\pi\sin(\vartheta)},\dots,\mathrm{e}^{-\mathrm{j}\pi(M-1)\sin(\vartheta)}\right]^{\mathrm{T}}.$ (7)
For every channel sample, we consider a new ${\bm{\delta}}$ and draw the sample according to ${\bm{h}}\sim\mathcal{N}_{\mathbb{C}}\left(\bm{0},{\bm{C}}_{\bm{\delta}}\right)$.
II-B Measurement Campaign
Since synthetic data capture real-world CSI characteristics only up to some extent, we utilize real-world data from a measurement campaign conducted at the Nokia campus in Stuttgart, Germany, in October/November 2017, cf. [36]. The BS antenna with a uniform rectangular array (URA) comprises $N_{v}=4$ vertical ($\lambda$ spacing) and $N_{h}=16$ horizontal ($\lambda/2$ spacing) single polarized patch antennas. The operating carrier frequency is $2.18$ GHz and the antenna was mounted on a rooftop approximately $20$ meters above the ground. For further details, we refer the reader to [36].
III Semi-Blind Channel Estimation using Perfect Statistical Knowledge
In this section, we introduce two variants of the LMMSE estimator incorporating information provided by the payload data symbols. To do so, we first restrict ourselves to the case where perfect statistical knowledge is available at the receiver.
In channel estimation, commonly, only the pilot observation ${\bm{y}}_{p}$ is considered for channel estimation. The MSE optimal estimator given the pilot observation ${\bm{y}}_{p}$ is the CME
$\hat{{\bm{h}}}_{\text{CME}}=\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p}\right].$ (8)
If the channel sample ${\bm{h}}$ is drawn from a Gaussian distribution according to
${\bm{h}}\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}),$ (9)
and if further this statistic is known at the receiver, the genie-aided CME can be formulated as [27, Ch. 10]
$\hat{{\bm{h}}}_{\text{CME}}={\bm{C}}\left({\bm{C}}+{\bm{C}}_{\bm{n}}\right)^{-1}{\bm{y}}_{p}.$ (10)
This LMMSE estimator achieves the MSE of
$\mathrm{MSE}^{\mathrm{plain}}=\mathrm{tr}\left[{\bm{C}}-{\bm{C}}({\bm{C}}+{\bm{C}}_{\bm{n}})^{-1}{\bm{C}}\right].$ (11)
For ${\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M}$, the MSE can be expressed using the Woodbury identity as
$\mathrm{MSE}^{\mathrm{plain}}=\sum_{i=1}^{M}\frac{{\rho}_{i}\sigma^{2}}{{\rho}_{i}+\sigma^{2}},$ (12)
where $\rho_{i}$ are the eigenvalues of ${\bm{C}}$.
For our considerations concerning semi-blind channel estimation, we assume knowledge about the subspace defined by $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$, where we denote with ${\bm{V}}$ the left singular vectors of ${\bm{H}}$, which span the same subspace as the columns of ${\bm{H}}$. Generally, we can formulate the ML estimate of ${\bm{H}}$ in view of (2) as [16]
$\min_{{\bm{H}},{\bm{X}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{h}}_{n}\right\|^{2}+\sum_{n=J+1}^{N}\left\|{\bm{y}}(n)-{\bm{H}}{\bm{x}}(n)\right\|^{2},$ (13)
where the first term belongs to the pilot observations and the second part refers to the observation obtained from the payload symbols. In [16] the EM algorithm is introduced to solve this channel estimation problem in terms of maximum likelihood, whereas, in [11] a semi-blind method is derived based on utilizing the subspace $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$. Now, given the subspace $\mathrm{range}({\bm{V}})$, we can reformulate (13) as [11]
$\min_{{\bm{S}},{\bm{X}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{s}}_{n}\right\|^{2}+\sum_{n=J+1}^{N}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{S}}{\bm{x}}(n)\right\|^{2},$ (14)
with ${\bm{H}}={\bm{V}}{\bm{S}}$ and ${\bm{S}}=[{\bm{s}}_{1},...,{\bm{s}}_{J}]\in\mathbb{C}^{J\times J}$. The optimal solution for ${\bm{X}}$ is given as
${\bm{x}}^{*}(n)$ $=({\bm{S}}^{\mathrm{H}}{\bm{V}}^{\mathrm{H}}{\bm{V}}{\bm{S}})^{-1}{\bm{S}}^{\mathrm{H}}{\bm{V}}^{\mathrm{H}}{\bm{y}}(n)$ (15)
$=({\bm{S}}^{\mathrm{H}}{\bm{S}})^{-1}{\bm{S}}^{\mathrm{H}}{\bm{V}}^{\mathrm{H}}{\bm{y}}(n).$ (16)
Reinserting this solution into (13), the right part of the objective simplifies to
$\min_{{\bm{S}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{s}}_{n}\right\|^{2}+\sum_{n=J+1}^{N}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}(n)\right\|^{2},$ (17)
where only the left part depends on ${\bm{S}}$. Thus, the ML problem for the user of interest results in [11]
$\min_{\bm{s}}\|{\bm{y}}_{p}-{\bm{V}}{\bm{s}}\|^{2},$ (18)
with the closed form solution $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$. The MSE of this estimator is
$\mathrm{MSE}^{\mathrm{ML}}=\mathbb{E}\left[\|{\bm{h}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}\|^{2}\right]=J\sigma^{2}.$ (19)
In the following, we introduce two channel estimation strategies combining the information provided by ${\bm{C}}$ and $\mathrm{range}({\bm{V}})$.
III-A Subspace Channel Estimator
Using the information in $\mathrm{range}({{\bm{V}}})$, we can solve the estimation within the subspace as previously proposed in [26]. For this, the pilot system model in (4) is projected into the $J$-dimensional subspace as
${\bm{y}}^{\prime}={\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$ $={\bm{V}}^{\mathrm{H}}{\bm{h}}+{\bm{V}}^{\mathrm{H}}{\bm{n}}={\bm{h}}^{\prime}+{\bm{n}}^{\prime}.$ (20)
Under the assumption that ${\bm{V}}$ is chosen independently of ${\bm{h}}$, the distribution ${\bm{h}}^{\prime}\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{V}}^{\mathrm{H}}{\bm{C}}{\bm{V}})$ can be used to formulate the estimate [26]
$\hat{{\bm{h}}}^{\prime}=\;$ ${\bm{V}}^{\mathrm{H}}{\bm{C}}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}.$ (21)
One should note that by design ${\bm{V}}$ actually depends on ${\bm{h}}$ and, hence, (21) is a suboptimal but feasible estimate for ${\bm{h}}^{\prime}$. After solving the estimation in the subspace for ${\bm{h}}^{\prime}$, the solution can be transformed back using [26]
$\hat{{\bm{h}}}_{\text{sub}}={\bm{V}}\hat{{\bm{h}}}^{\prime}.$ (22)
III-B Projected Channel Estimator
As an alternative approach, we propose the use of the orthogonal subspace projection ${\bm{P}}_{\bm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ as a preprocessing filter. Since the projector ${\bm{P}}_{\bm{H}}$ does not affect ${\bm{h}}$, the resulting projected observation is given by
$\tilde{{\bm{y}}}={\bm{P}}_{\bm{H}}{\bm{y}}_{p}={\bm{h}}+{\bm{P}}_{\bm{H}}{\bm{n}}={\bm{h}}+\tilde{{\bm{n}}}.$ (23)
To formulate the CME $\hat{{\bm{h}}}=\mathbb{E}\left[{\bm{h}}\mid\tilde{{\bm{y}}}\right]$, we need to calculate the statistic of the noise $\tilde{{\bm{n}}}$ with
${\bm{C}}_{\tilde{{\bm{n}}}}=\;$ $\mathbb{E}\left[\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right]=\mathbb{E}\left[\sigma^{2}{\bm{P}}_{\bm{H}}\right].$ (24)
To get an intuitive understanding of (24), let us consider a scenario involving spatially uncorrelated channels, meaning that path gains and channel directions are uncorrelated. This is the case when users are uniformly distributed over the directions, e.g., the spatial channel model in Section II-A, resulting in a channel covariance matrix of the scenario that is a scaled identity [37, Def. 2.3]. In such a case, the matrices with the eigenvectors of the sample covariance matrix of such channels are distributed with Haar measure [38, Chap. 1], i.e., uniformly distributed on the manifold of unitary matrices. Assuming spatially uncorrelated channels (24) results in
${\bm{C}}_{\tilde{{\bm{n}}}}=\sigma^{2}\frac{J}{M}\mathbf{I}_{M},$ (25)
which we assume to hold for the remainder of this section. We can then formulate the projected LMMSE estimator as
$\hat{{\bm{h}}}_{\text{proj}}={\bm{C}}\left({\bm{C}}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}\tilde{{\bm{y}}}.$ (26)
III-C Performance Analysis
If (25) is true, the MSE of the proposed projected channel estimator can directly be written as (cf. Section -A)
$\mathrm{MSE}^{\mathrm{proj}}$ $=\mathrm{tr}\left({{\bm{C}}}-{{\bm{C}}}\left({{\bm{C}}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{{\bm{C}}}\right)$ (27)
$=\sum_{i=1}^{M}\frac{\rho_{i}\sigma^{2}}{\frac{M}{J}\rho_{i}+\sigma^{2}}.$ (28)
Comparing the performance to the plain LMMSE we see that
$\frac{\rho_{i}\sigma^{2}}{\frac{M}{J}\rho_{i}+\sigma^{2}}\leq\frac{\rho_{i}\sigma^{2}}{\rho_{i}+\sigma^{2}},$ (29)
holds for every $i=1,\dots,M$, resulting in $\mathrm{MSE}^{\mathrm{proj}}\leq\mathrm{MSE}^{\mathrm{plain}}$, cf. (12). The inequality in (29) only holds with equality if $J=M$. Additionally, we can compare the MSE of the projected LMMSE to (19) by reformulating (28) as
$\mathrm{MSE}^{\mathrm{proj}}=\frac{J}{M}\sigma^{2}\sum_{i=1}^{M}\frac{\rho_{i}}{\rho_{i}+\frac{J}{M}\sigma^{2}}\leq J\sigma^{2}=\mathrm{MSE}^{\text{ML}}.$ (30)
To compare the projected variant to the subspace LMMSE, let us consider uncorrelated Rayleigh fading such that the channel covariance matrix is given as ${\bm{C}}=\mathbf{I}_{M}$. The MSE of the projected variant results in
$\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}=\frac{JM\sigma^{2}}{{M}+J\sigma^{2}}.$ (31)
For the case of ${\bm{C}}=\mathbf{I}_{M}$, the subspace LMMSE estimator boils down to
$\hat{{\bm{h}}}_{\mathrm{sub}}=\frac{1}{1+\sigma^{2}}{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}},$ (32)
with its corresponding MSE as (cf. Section -B)
$\mathrm{MSE}^{\mathrm{sub}}_{\mathrm{iid}}=\frac{\sigma^{2}(M\sigma^{2}+J)}{(1+\sigma^{2})^{2}}\geq\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}.$ (33)
Fig. 1 shows the performance of the individual channel estimators based on perfect genie knowledge for the case of uncorrelated Rayleigh fading. As one can see, the projected channel estimator outperforms all other estimators across the whole signal-to-noise ratio (SNR) range. Further, we realize that the subspace LMMSE converges to the ML method from above.
Figure 1: MSE over the SNR for given channel estimations based on perfect subspace and perfect statistical knowledge in a $J=8$ user and $M=64$ antennas scenario with uncorrelated Rayleigh fading.
IV Proposed Semi-Blind Channel Estimation - Utilizing Generative Prior
In practice, (10) can not be utilized directly as the channels have to be Gaussian distributed, and this distribution needs to be known. In general, the CME given the pilot observation ${\bm{y}}_{p}$ is formulated as
$\hat{{\bm{h}}}_{\text{CME}}=\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p}\right]=\int{\bm{h}}\frac{p_{{\bm{n}}}({\bm{y}}_{p}-{\bm{h}})p({\bm{h}})}{p({\bm{y}}_{p})}\mathrm{d}{\bm{h}}.$ (34)
As can be seen in (34), the CME generally can not be computed analytically. First, the CME needs access to $p({\bm{h}})$, which is generally unavailable in practice. Additionally, no closed-form solution exists to the integral in (34).
In order to reformulate the CME, we first use the property that for any arbitrarily distributed random variable ${\bm{h}}$, we can always find a condition ${\bm{c}}$ which makes the conditional distribution Gaussian. Secondly, it has been shown in [39] that for wireless communication channels, this conditional Gaussian distribution preserves the zero-mean property as
${\bm{h}}\mid{\bm{\delta}}\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{{\bm{h}}\mid{\bm{c}}}).$ (35)
Thus, we can reformulate the CME as
$\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p},{\bm{c}}\right]\mid{\bm{y}}_{p}\right]$ (36)
$=\int\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p},{\bm{c}}\right]p({\bm{c}}\mid{\bm{y}}_{p})\mathrm{d}{\bm{c}}$ (37)
$\approx\int\hat{{\bm{h}}}_{\bm{c}}({\bm{y}}_{p})p({\bm{c}}\mid{\bm{y}}_{p})\mathrm{d}{\bm{c}}$ (38)
where
$\hat{{\bm{h}}}_{\bm{c}}({\bm{y}}_{p})={\bm{C}}_{{\bm{h}}\mid{\bm{c}}}\left({\bm{C}}_{{\bm{h}}\mid{\bm{c}}}+{\bm{C}}_{\bm{n}}\right)^{-1}{\bm{y}}_{p},$ (39)
denotes the LMMSE estimate given ${\bm{c}}$. However, finding a suitable condition $\bm{\delta}$ can be challenging, in particular as the true distribution of ${\bm{h}}$ is unknown. To this end, CGLMs were proposed in [30, 31, 32], which approximate the CME based on a GMM, mixtures of factor analyzers (MFA), and VAE, respectively. All three methods learn a model that provides the conditional Gaussian distribution ${\bm{h}}\mid{\bm{c}}$ based on a discrete (GMM, MFA) or continuous (VAE) latent variable ${\bm{c}}$. In this work, we focus on the GMM and VAE, which we adapt to semi-blind channel estimation in the following.
IV-A GMM-based Semi-blind Channel Estimation
Based on the universal approximation property of GMMs [40], the PDF of ${\bm{h}}$ is approximated by
$f_{\bm{h}}^{(K)}({\bm{h}})=\sum_{k=1}^{K}p(k)\mathcal{N}_{\mathbb{C}}({\bm{h}};\bm{\mu}_{k},{\bm{C}}_{k}),$ (40)
where $p(k)$, $\bm{\mu}_{k}$ and ${\bm{C}}_{k}$ are the mixing coefficients, means and covariance matrices of the $k$-th GMM component, respectively. As we are considering wireless channels, the mean of each component is set to $\bm{\mu}_{k}=\bm{0}$, cf. [39]. The fitting of the components in (40) is accomplished with the well-known EM algorithm [41] based on a set =$\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ of $T$ channel samples as training data. Based on the formulation in (40) the conditional PDF given $k$ is
${\bm{h}}\mid k\sim\mathcal{N}_{\mathbb{C}}({\bm{h}};\bm{0},{\bm{C}}_{k}).$ (41)
Thus, in the case of a GMM we have a discrete latent variable, which helps us parameterize the CME. The resulting semi-blind subspace GMM can be formulated as
$\hat{{\bm{h}}}_{\text{sub. GMM}}={\bm{V}}\hat{{\bm{h}}}^{\prime}_{\text{GMM}}={\bm{V}}\sum_{k=1}^{K}p(k\mid{\bm{y}}^{\prime})\hat{{\bm{h}}}^{\prime}_{\text{GMM},k},$ (42)
with
$\hat{{\bm{h}}}^{\prime}_{\text{GMM},k}=\;$ ${\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p},$ (43)
and the corresponding responsibilities
$p(k\mid{\bm{y}}^{\prime})=\frac{p(k)\mathcal{N}_{\mathbb{C}}\left({\bm{y}}^{\prime};\bm{0},{\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)}{\sum_{i=1}^{K}p(i)\mathcal{N}_{\mathbb{C}}\left({\bm{y}}^{\prime};\bm{0},{\bm{V}}^{\mathrm{H}}{\bm{C}}_{i}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)}.$ (44)
The projected GMM is
$\hat{{\bm{h}}}_{\text{proj. GMM}}=\sum_{k=1}^{K}p(k\mid\tilde{{\bm{y}}})\hat{{\bm{h}}}_{\text{proj. GMM},k},$ (45)
with
$\hat{{\bm{h}}}_{\text{proj. GMM},k}={\bm{C}}_{k}\left({\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}\tilde{{\bm{y}}}$ (46)
and the associated responsibilities
$p(k\mid\tilde{{\bm{y}}})=\frac{p(k)\mathcal{N}_{\mathbb{C}}\left(\tilde{{\bm{y}}};\bm{0},{\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)}{\sum_{i=1}^{K}p(i)\mathcal{N}_{\mathbb{C}}\left(\tilde{{\bm{y}}};\bm{0},{\bm{C}}_{i}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)}.$ (47)
The respective estimators are summarized in Algorithm 1 and Algorithm 2.
Algorithm 1 Subspace GMM Channel Estimator
Offline Training Phase
1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the GMM with the EM algorithm, cf. [30]
3:
4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
8:for $j=1,\dots,J$ do
9:     ${\bm{y}}^{\prime}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:     for $k=1,\dots,K$ do
11:         $\hat{{\bm{h}}}^{\prime}_{k}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{y}}^{\prime}$
12:     end for
13:     $\hat{{\bm{h}}}_{j}\leftarrow{\bm{V}}\sum_{k=1}^{K}p(k\mid{\bm{y}}^{\prime})\hat{{\bm{h}}}^{\prime}_{k}$
14:end for
15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
Algorithm 2 Projected GMM Channel Estimator
Offline Training Phase
1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the GMM with the EM algorithm, cf. [30]
3:
4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
8:for $j=1,\dots,J$ do
9:     $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:     for $k=1,\dots,K$ do
11:         $\hat{{\bm{h}}}_{k}\leftarrow{\bm{C}}_{k}\left({\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}\tilde{{\bm{y}}}$
12:     end for
13:     $\hat{{\bm{h}}}_{j}\leftarrow\sum_{k=1}^{K}p(k\mid\tilde{{\bm{y}}})\hat{{\bm{h}}}_{k}$
14:end for
15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
Figure 2: Structure of a semi-blind VAE. The matrix $\hat{{\bm{V}}}$ contains the $J$ dominant eigenvectors of (57). The encoder and decoder represent DNNs.
IV-B VAE-based Semi-blind Channel Estimation
To learn the unknown distribution $f_{\bm{h}}({\bm{h}})$ using a VAE, we lower bound the parameterized likelihood $p_{\bm{\theta}}({\bm{h}})$ using the evidence-lower bound (ELBO). To formulate the ELBO, the variational distributions $q_{\bm{\phi}}({\bm{z}}\mid{\bm{y}}^{\prime})$ and $q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$ are introduced, which approximate $p({\bm{z}}\mid{\bm{y}}^{\prime})$ and $p({\bm{z}}\mid\tilde{{\bm{y}}})$, respectively. In contrast to the GMM, the used subspace $\mathrm{range}({\bm{V}})$ is unknown to the encoder of the VAE making $p({\bm{z}}\mid{\bm{y}}^{\prime})$ difficult to learn. Additionally, the dimension of the encoder input would depend on the number of users in the system. Thus, we propose to approximate both posteriors $p({\bm{z}}\mid{\bm{y}}^{\prime})$ and $p({\bm{z}}\mid\tilde{{\bm{y}}})$ with $q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$. A version of the ELBO for this case, which is accessible, can be written as [42]
$\mathcal{L}_{{\bm{\theta}},{\bm{\phi}}}=\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}})]-\mathrm{D}_{\mathrm{KL}}(q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})\mid\mid p({\bm{z}})),$ (48)
where $\mathbb{E}_{q_{\bm{\phi}}}[\cdot]=\mathbb{E}_{q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})}[\cdot]$ is the expectation over the variational distribution $q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$. The second term in (48) is the Kullback-Leibler (KL) divergence
$\mathrm{D}_{\mathrm{KL}}(q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})\mid\mid p({\bm{z}}))=\mathbb{E}_{q_{\bm{\phi}}}\left[\log\left(\frac{q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})}{p({\bm{z}})}\right)\right].$ (49)
In the VAE framework, the ELBO is optimized using deep neural networks (DNNs) and the reparameterization trick [42]. In order to do so, the involved distributions are defined as
$p({\bm{z}})$ $=\mathcal{N}(\bm{0},\mathbf{I}_{Z}),$
$p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}})$ $=\mathcal{N}_{\mathbb{C}}(\bm{\mu}_{\bm{\theta}}({\bm{z}}),{{\bm{C}}}_{\bm{\theta}}({\bm{z}})),$ (50)
$q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$ $=\mathcal{N}(\bm{\mu}_{\bm{\phi}}(\tilde{{\bm{y}}}),\mathrm{diag}(\bm{\sigma}^{2}_{\bm{\phi}}(\tilde{{\bm{y}}}))).$
The resulting semi-blind VAE structure is shown in Fig. 2. In the case of a ULA or URA at the BS the channel covariance matrix is either Toeplitz or block-Toeplitz, respectively. As shown in [39], the conditional covariance matrix at the output of the VAE preserves this structure. Thus, we parameterize the output covariance matrix as
${{\bm{C}}}_{\bm{\theta}}({\bm{z}})={\bm{Q}}^{\mathrm{H}}\mathrm{diag}({\bm{c}}_{\bm{\theta}}({\bm{z}})){\bm{Q}},$ (51)
where ${\bm{Q}}={\bm{Q}}_{M}$ or ${\bm{Q}}={\bm{Q}}^{\prime}_{N_{v}}\otimes{\bm{Q}}^{\prime}_{N_{h}}$, respectively, where ${\bm{Q}}_{M}$ is a DFT matrix of size $M$ resulting in a circulant approximation, cf. [32], and ${\bm{Q}}^{\prime}_{N_{x}}$ contains the first $N_{x}$ columns of the $2N_{x}\times 2N_{x}$ DFT matrix resulting in a block-Toeplitz parameterization, cf. [43]. Further, for the (block-)Toeplitz parameterization we can set $\bm{\mu}_{\bm{\theta}}({\bm{z}})=\bm{0}$, cf. [39].
After successfully training the VAE, the output is a local parameterization of $f_{\bm{h}}({\bm{h}})$ as conditionally Gaussian
${\bm{h}}\mid{\bm{z}}\sim p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}}).$ (52)
As analyzed in [32] it is a reasonable approximation to set
$p({\bm{z}}\mid\tilde{{\bm{y}}})=\begin{cases}1\quad\text{if }{\bm{z}}=\bm{\mu}_{\bm{\phi}}(\tilde{{\bm{y}}}),\\ 0\quad\text{otherwise}.\end{cases}$ (53)
Based on this parameterization we can formulate the semi-blind VAE-based estimators as
[…]
and
[…]
×$\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}).$ (55)
The respective estimators are summarized in Algorithm 3 and Algorithm 4. For a more detailed introduction into the VAE framework and its usage for parameterization of the CME we refer the reader to [32].
Algorithm 3 Subspace VAE Channel Estimator
Offline Training Phase
1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the VAE by optimizing the ELBO, cf. [32]
3:
4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
8:for $j=1,\dots,J$ do
9:     $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:     $\bm{\mu}_{\bm{\theta}}({\bm{z}}),{\bm{C}}_{\bm{\theta}}({\bm{z}})\leftarrow\mathrm{VAE}(\tilde{{\bm{y}}})$
11:     $\hat{{\bm{h}}}_{j}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}$
12:           ×$\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}})$
13:end for
14:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
Algorithm 4 Projected VAE Channel Estimator
Offline Training Phase
1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the VAE by optimizing the ELBO, cf. [32]
3:
4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
8:for $j=1,\dots,J$ do
9:     $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:     $\bm{\mu}_{\bm{\theta}}({\bm{z}}),{\bm{C}}_{\bm{\theta}}({\bm{z}})\leftarrow\mathrm{VAE}(\tilde{{\bm{y}}})$
11:     $\hat{{\bm{h}}}_{j}\leftarrow\bm{\mu}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\bm{\theta}}({\bm{z}})\left({\bm{C}}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}}))$
12:           ×$\times(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}}))$
13:end for
14:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
IV-C Maximum Likelihood Subspace Estimation
After introducing the methods utilizing the additional subspace information provided by $\mathrm{range}({\bm{V}})$ to enhance the CSI estimation quality, let us consider the estimation of such a subspace. As the received data symbols are also transmitted over the same channel, they can be used to estimate the subspace containing all user channels.
To this end, let us reconsider the ML estimate of ${\bm{H}}$ in (13). Instead of directly optimizing on this ML formulation as done in [16], which generally does not result in the MMSE, we only take this log-likelihood formulation as an intermediate step to estimate the subspace $\mathrm{range}({\bm{V}})$. First, let us consider the right term of the objective function in (13). We can then reformulate the problem by again solving for ${\bm{X}}$ first and reinserting the solution resulting in [11]
$\max_{\bm{H}}\;\mathrm{tr}\left({\bm{P}}_{\bm{H}}\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}\right),$ (56)
where ${\bm{P}}_{\bm{H}}={\bm{H}}({\bm{H}}^{\mathrm{H}}{\bm{H}})^{-1}{\bm{H}}^{\mathrm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ and $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N-J}{\bm{Y}}_{d}{\bm{Y}}_{d}^{\mathrm{H}}$, with ${\bm{Y}}_{d}$ from (2). The maximization in (56) is solved by setting ${\bm{P}}_{\bm{H}}$ equal to $\hat{{\bm{V}}}_{d}\hat{{\bm{V}}}_{d}^{\mathrm{H}}$ with $\hat{{\bm{V}}}_{d}$ holding the $J$ dominant eigenvectors of the receive sample covariance matrix $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}$. This result has also been used in [26]. Additionally, it is trivial to see that the first term in (13) is minimized by ${\bm{h}}_{n}={\bm{y}}(n)$. The subspace spanned by the solution ${\bm{h}}_{n}={\bm{y}}(n)$ is the same as the subspace spanned by the $J$ eigenvectors of the sample covariance matrix $\hat{{\bm{C}}}^{(p)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{J}{\bm{Y}}_{p}{\bm{Y}}_{p}^{\mathrm{H}}$, which ignores the additional phase information contained in the pilot observation. Thus, the overall subspace estimate $\hat{{\bm{V}}}=[{\bm{v}}_{1},\dots,{\bm{v}}_{J}]$ is found by taking the $J$ dominant eigenvectors ${\bm{v}}_{j}$ of the sample covariance matrix defined as
$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}.$ (57)
To utilize information from the previous coherence intervals, one can adaptively update the subspace using efficient tracking algorithms as proposed in, e.g., [44, 45].
IV-D Complexity Analysis
The standalone GMM estimator proposed by [30] precomputes the filters used for the individual components, resulting in a complexity of $\mathcal{O}(KM^{2})$. For the standalone VAE the complexity is given as $\mathcal{O}(DM^{2})$ [32], where $D$ denotes the number of layers in the forward pass of the VAE. For our semi-blind methods, the calculation of the subspace requires $\mathcal{O}((N+J)M^{2})$. This results from calculating the sample covariance matrix with $\mathcal{O}(NM^{2})$ and taking the eigenvectors of the $J$ largest eigenvalues for the solution of (56). Using the projection approximation subspace tracking (PAST) algorithm [45], the computational complexity of calculating the subspace reduces to $\mathcal{O}(JM)$ for every update. In the case of the subspace GMM the $K$ LMMSE estimates can not be precomputed, which results in a complexity of $\mathcal{O}(K(M^{2}+JM^{2}+J^{3}))$. Similarily, the subspace VAE exhibits a complexity of $\mathcal{O}(DM^{2}+JM^{2}+J^{3})$. For the projected versions of the GMM and VAE the complexity becomes $\mathcal{O}(KM^{2}+JM^{2})$ and $\mathcal{O}(DM^{2}+JM^{2})$, respectively. One should note that the calculation for each of the $K$ components in the GMM can be parallelized. Similarly, the computations in the convolutional layers of the VAE can be parallelized, mitigating the complexity.
V Baseline estimators
To compare our methods, the following baseline channel estimators are considered. Based on the found subspace $\mathrm{range}({\bm{V}})$ we can formulate the pilot-based ML estimator as $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$, which is the closed-form solution to (18). This can be interpreted as the subspace-adjusted version of the conventional least squares (LS) channel estimator given as $\hat{{\bm{h}}}_{\text{LS}}={\bm{y}}_{p}$.
Another estimator is based on the sample covariance matrix, which we can compute from the training data set ${\mathcal{H}}$ to infer the global statistics of the channels as
${\bm{C}}_{s}=\frac{1}{|{\mathcal{H}}|}\sum_{{{\bm{h}}}\in{\mathcal{H}}}{{\bm{h}}}{{\bm{h}}}^{\mathrm{H}}.$ (58)
We can use the matrix ${\bm{C}}$ as statistical prior to parameterize the semi-blind channel estimators outlined in Section III-A and Section III-B as
$\hat{{\bm{h}}}_{\text{sub. s-cov}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{C}}_{s}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{s}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p},$ (59)
and
$\hat{{\bm{h}}}_{\text{proj. s-cov}}={\bm{C}}_{s}\left({\bm{C}}_{s}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}{\bm{P}}_{\bm{H}}{\bm{y}}_{p}.$ (60)
Lastly, we compare our proposed methods to two iterative algorithms optimizing the ML formulation in (13), namely the EM from [16] and a MP variant similar to [17], which we run both until convergence or $500$ iteration, whatever comes first.
VI Numerical Simulations
(a) Spatial Channel Model (Sec. II-A)
(b) Measurement Data (Sec. II-B)
Figure 3: NMSE over the SNR for given channel estimations based on $N=200$ observations including one pilot per user in a $J=8$ user scenario using (a) the spatial channel model and (b) measurement data.
To evaluate our proposed methods, we use channel realizations, which are normalized with $\mathbb{E}\left[\|{\bm{h}}\|^{2}\right]=M$. Thus, we can define the SNR $=\frac{1}{\sigma^{2}}$. Further, the normalized MSE (NMSE) defined as
$\text{NMSE}=\frac{1}{ML}\sum_{\ell=1}^{L}\|{\bm{h}}_{\ell}-\hat{{\bm{h}}}_{\ell}\|^{2},$ (61)
is used to characterize the performance of the estimators based on $L=10^{3}$ unseen channel samples stemming from the channel models detailed in Sections II-A and II-B. The assumption of spatial uncorrelated channels, as used in (25), only holds for the spatial channel model of Section II-A. For the case of the measurement campaign described in Section II-B, we approximate the noise covariance matrix (24) as
${\bm{C}}_{\tilde{{\bm{n}}}}\approx\sigma^{2}\frac{J}{M}\mathbf{I}_{M}.$ (62)
We use =$\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ with $T=1.5\cdot 10^{5}$ training samples from the respective channel model to train the GMM and VAE, where we set the number of components to $K=64$ and the latent dimension to $Z=32$, respectively. Further, in the case of the VAE we allow non-zero values for $\bm{\mu}_{\bm{\theta}}({\bm{z}})$, as we use the circulant approximation for the spatial channel model and the block-Toeplitz property is not perfectly fulfilled for the measurement data, due to hardware imperfections. The “s-cov” variants (“sub. s-cov” and “proj. s-cov”) utilize the same training samples. The number of BS antenna is set to $M=64$, cf. Sections II-A and II-B, serving $J=8=M/8$ number of users, a representative operating point [37, Chap. 1.3.3]. Further, the number of snapshots is set to $N=200$, if not stated otherwise, corresponding to a scenario that allows high channel dispersion and high mobility, e.g., up to $135$ kph, c.f. [37, Chap. 2.1]. The sent symbols during data transmission generally stem from a discrete constellation, e.g., QPSK, $16$-QAM. For this work, we utilize Gaussian symbols with $x_{j}(n)\sim\mathcal{N}_{\mathbb{C}}(0,P_{j}=1/J)$ such that $\sum_{j=1}^{J}P_{j}=1$. Using a continuous symbol constellation has a negligible effect on the results of the simulations, as also previously observed in [16].
Fig. 3a and Fig. 3b show the performance of the different channel estimation methods with respect to the SNR for the spatial channel model (cf. Section II-A) and measurement data (cf. Section II-B), respectively. One can see that the semi-blind methods utilizing the CGLMs perform the best across the whole SNR. The projected variant slightly outperforms its subspace counterpart for most SNR values, which follows the derivations in Section III. Interestingly, the order of the semi-blind GMM and semi-blind VAE depends on the utilized channel model. In Fig. 3a, the projected GMM and VAE show both the best overall result, whereas in Fig. 3b, the projected GMM outperforms all other estimators. Additionally, we see that in Fig. 3a, the subspace VAE outperforms the subspace GMM and in Fig. 3b, the results are vice versa. This ordering follows the ordering of the standalone version, where the plain GMM is better than the plain VAE in the case of the measurement data and worse for the spatial channel model. For high SNR values, all semi-blind variants approach each other except the EM and MP methods. In Fig. 3a, the EM and MP drastically improve from $15$dB to $20$dB showing similar performance at $20$dB as the other semi-blind methods, whereas, in Fig. 3b, they show inferior results also for high SNR. The CGLM-based approaches keep a slight advantage even for high SNR values, which can be attributed to the fact that prior information is beneficial even for high SNR. A notable observation is that, in the mid-SNR range, the semi-blind CGLM variants outperform all related estimators by roughly $3$ dB.
(a) Spatial Channel Model (Sec. II-A)
(b) Measurement Data (Sec. II-B)
Figure 4: NMSE over the number of observations for given channel estimations based on $N$ observations including one pilot per user at SNR $=0$ dB in a $J=8$ user scenario using (a) the spatial channel model and (b) measurement data.
For our proposed strategies, the accuracy of the estimated subspace $\mathrm{range}(\hat{{\bm{V}}})$ influences the performance and, hence, the NMSE depends on the number of snapshots $N$ as shown in Fig. 4. We see that for an increasing number of snapshots, the NMSE of our proposed methods decreases. In the case of the spatial channel model (Fig. 4a) the projected GMM and VAE perform best for low numbers of snapshots, where the standalone VAE surpasses all other methods for $N=20$. Additionally, we observe in Fig. 4a that for high $N$, the subspace VAE becomes the best of all considered methods. In Fig. 4b, we observe again that for the measurement data, the semi-blind GMM variants perform the best, where for high $N$ the subspace GMM and low $N$ the projected GMM outperforms all other methods. Again, for less than $30$ snapshots, the semi-blind methods are outperformed by the standalone GMM and VAE due to inaccuracies in estimating the subspace with a low number of payload data symbols. For both utilized channel models, the subspace variant of the superior CGLM surpasses the projected variant for high numbers of snapshots converging to a lower error level. Thus, in practice, where, in general, uncorrelated Rayleigh fading is not the case, there are cases where the subspace CGLM outperforms its projected counterpart.
A critical decrease in performance can be observed for the EM and MP algorithms. Here, the NMSE increases after a certain point when increasing the number of snapshots. Even though the minimum appears at different $N$, the overall behavior exhibits similarities. This is because both methods optimize the joint ML formulation in (13), where the optimization of the second term becomes dominant for a high number of snapshots. Hence, the impact of the pilot observations relevant to estimating the phase of the channel vanishes.
(a) Spatial Channel Model (Sec. II-A)
(b) Measurement Data (Sec. II-B)
Figure 5: NMSE over the number of users for given channel estimations based on $N=200$ observations including one pilot per user at SNR $=0$ dB using (a) the spatial channel model and (b) measurement data.
The dimension of the subspace $\mathrm{range}(\hat{{\bm{V}}})$ directly influences the estimation quality of the proposed methods as shown in Fig. 5a and Fig. 5b. For example, in the extreme case where the number of users in the system is equal to the number of BS antennas ($J=M$), the solution to (56) becomes $\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}=\mathbf{I}$ and, hence, as the number of users in the system increases all semi-blind estimators approach their respective purely pilot based version. We restrict our simulations within the interval of $J\leq M/4=16$, which is said to be the preferred operating regime in massive MIMO [37, Chap. 1.3.3], and set the number of snapshots to $N=200$. In the case of a single user, all semi-blind variants exhibit similar performance, except for the subspace sample covariance estimator, the subspace VAE, and in the case of the spatial channel model (Fig. 5a) the subspace GMM. For all other considered numbers of users, the proposed projected CGLM methods outperform all other channel estimators. Additionally, for the spatial channel model in Fig. 5a, the subspace GMM also shows inferior results to the other CGLM-based methods for all numbers of users.
Overall, we can conclude that the proposed semi-blind CGLMs show superior channel estimation performance across all different setups. Depending on the used channel model, either the semi-blind GMMs or the semi-blind VAEs result in slightly better NMSE, where only in the case of the spatial channel model, the subspace GMM shows slightly worse performance compared to the other proposed methods. Moreover, the projected CGLMs outperform their respective subspace counterpart for most simulated operating points, showing the superiority of the proposed projection method.
VII Conclusion
This work presented a novel semi-blind channel estimation technique based on the class of CGLMs. To this end, two methods are discussed that incorporate subspace knowledge about the channel into the well-known LMMSE estimator. Both methods exploit the estimated subspace derived from the dominant eigenvectors of sample covariance matrices constructed using the received symbols. A theoretical analysis of the methods showed the superior estimation quality of the proposed projection-based estimator for uncorrelated Rayleigh fading channels. Secondly, we showed how two examples from the class of CGLMs, i.e., the GMM and VAE, can be used to parameterize these estimators. Extensive simulations based on real-world measurement and spatial channel model data demonstrated the superior estimation performance of the proposed methods compared to standard semi-blind channel estimators.
-A MSE of Projected LMMSE
For any linear estimator $\hat{{\bm{h}}}={\bm{W}}{\bm{y}}_{p}$, the MSE is given as
$\mathrm{MSE}$ $=\mathbb{E}\left[\|{\bm{h}}-\hat{{\bm{h}}}\|^{2}\right]=\mathbb{E}\left[\mathrm{tr}\left(\left({\bm{h}}-\hat{{\bm{h}}}\right)\left({\bm{h}}^{\mathrm{H}}-\hat{{\bm{h}}}^{\mathrm{H}}\right)\right)\right]$ (63)
$=\mathbb{E}\left[\mathrm{tr}({\bm{h}}{\bm{h}}^{\mathrm{H}})-2\mathrm{tr}({\bm{h}}{\bm{y}}^{\mathrm{H}}{\bm{W}})+\mathrm{tr}({\bm{W}}{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}})\right].$ (64)
For the case of the projected LMMSE the second term in (64) can be rewritten as
$\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}\tilde{{\bm{y}}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}}\right)\right]$ (65)
$\quad\quad=\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}}\right)\right]$ (66)
$\quad\quad=\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\right)\right]$ (67)
$\quad\quad=\mathrm{tr}\left({\bm{C}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\right).$ (68)
Similarly for the third term in (64) we have
$\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}\tilde{{\bm{y}}}\tilde{{\bm{y}}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}}\right)\right]$ (69)
$\quad\quad=\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}+\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right){\bm{W}}^{\mathrm{H}}\right)\right]$ (70)
$\quad\quad=\mathbb{E}\bigg{[}\mathrm{tr}\bigg{(}{\bm{C}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}+\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right)$
×$\quad\quad\quad\times\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\bigg{)}\bigg{]}$ (71)
$\quad\quad=\mathrm{tr}\left({\bm{C}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\right),$ (72)
where we assume that $\mathbb{E}\left[\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right]=\sigma^{2}\frac{J}{M}\mathbf{I}_{M}$. From this the overall MSE in (28) follows directly.
-B MSE of Subspace LMMSE for Rayleigh Fading
In the case of uncorrelated Rayleigh fading the subspace LMMSE filter is given as
${\bm{W}}_{\mathrm{sub}}=\frac{1}{1+\sigma^{2}}{\bm{V}}{\bm{V}}^{\mathrm{H}}.$ (73)
Using this filter the second term in (64) can be rewritten as
$\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{y}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\right]$ (74)
$\quad\quad=\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$ (75)
$\quad\quad=\frac{1}{1+\sigma^{2}}\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{V}}{\bm{V}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$ (76)
$\quad\quad=\frac{1}{1+\sigma^{2}}\mathbb{E}_{\bm{h}}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}\right)\right]$ (77)
$\quad\quad=\frac{1}{1+\sigma^{2}}M.$ (78)
Similarly for the third term in (64) we have
$\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}_{\mathrm{sub}}{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\right]$ (79)
$\quad\quad=\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}_{\mathrm{sub}}{\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$
$\quad\quad\quad+\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}_{\mathrm{sub}}{\bm{n}}{\bm{n}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$ (80)
$\quad\quad=\frac{1}{(1+\sigma^{2})^{2}}\Big{[}\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{V}}{\bm{V}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$
$\quad\quad\quad+\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{n}}{\bm{n}}^{\mathrm{H}}{\bm{V}}{\bm{V}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]\Big{]}$ (81)
$\quad\quad=\frac{1}{(1+\sigma^{2})^{2}}\left[\mathbb{E}_{\bm{h}}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}\right)\right]+\mathbb{E}_{\bm{h}}\left[\sigma^{2}\mathrm{tr}\left({\bm{V}}{\bm{V}}^{\mathrm{H}}\right)\right]\right]$ (82)
$\quad\quad=\frac{1}{(1+\sigma^{2})^{2}}(M+J\sigma^{2}).$ (83)
The overall MSE of the subspace variant for ${\bm{C}}=\mathbf{I}_{M}$ is then
$\mathrm{MSE}_{\mathrm{iid}}^{\mathrm{sub}}$ $=M-2\frac{1}{1+\sigma^{2}}M+\frac{1}{(1+\sigma^{2})^{2}}(M+J\sigma^{2})$ (84)
$=\frac{\sigma^{2}(M\sigma^{2}+J)}{(1+\sigma^{2})^{2}}.$ (85)

### v2

@@
 To this end, we first estimate a subspace based on all received symbols (pilot and payload) to provide additional information for subsequent channel estimation.
-We show how this additional information enhances minimum mean square error (MMSE) channel estimation.
+This additional information enhances minimum mean square error (MMSE) channel estimation.
 Two variants of the linear MMSE (LMMSE) estimator are formulated, where the first one solves the estimation within the subspace, and the second one uses a subspace projection as a preprocessing step.
-Theoretical derivations show the superior estimation performance of the latter method in terms of mean square error for uncorrelated Rayleigh fading.
-Subsequently, we introduce parameterizations of this semi-blind LMMSE estimator based on two different conditional Gaussian latent models, i.e., the Gaussian mixture model and the variational autoencoder.
-Both models learn the underlying channel distribution of the propagation environment based on training data and serve as generative priors for semi-blind channel estimation.
-Extensive simulations for real-world measurement data and spatial channel models show the superior performance of the proposed methods compared to state-of-the-art semi-blind channel estimators with respect to the MSE.
-Index Terms:
-Semi-blind channel estimation, Gaussian mixture model, variational autoencoder, measurement data.
+Theoretical derivations show the latter method’s superior estimation performance in terms of mean square error for uncorrelated Rayleigh fading.
+Further, we provide asymptotical insights on how the proposed MMSE-based channel estimation strategy outperforms the unbiased Cramer-Rao bound.
+Subsequently, we introduce parameterizations of these semi-blind LMMSE estimators based on two different conditional Gaussian latent models, i.e., the Gaussian mixture model and the variational autoencoder.
+Both models learn the propagation environment’s underlying channel distribution based on training data and serve as generative priors for our semi-blind channel estimation.
+Extensive simulations for real-world measurement data and spatial channel models show the proposed methods’ superior performance compared to state-of-the-art semi-blind channel estimators in terms of MSE.
 I Introduction
 Accurate channel state information (CSI) is crucial for achieving the expected high data rates promised by multiple-input-multiple-output (MIMO) systems [2, 3, 4].
-The CSI describes the communication link between transmitter and receiver, characterized by its time-varying and frequency-selective nature, which is prone to rapid changes making the task of channel estimation complex [5].
-As accurate channel estimates are essential for the successful transmission of data, it is placed at the center of several research efforts [6, 7].
-The most widely adopted methods in wireless communication utilize known training or pilot symbols transmitted across the channel using some of the radio resource blocks [8].
+The CSI describes the communication link between transmitter and receiver, characterized by its time-varying and frequency-selective nature, which is prone to rapid changes, making the task of channel estimation complex [5].
+As accurate channel estimates are essential for successfully transmitting data, it is at the center of several research efforts [6, 7].
+The most widely adopted methods in wireless communications utilize known training or pilot symbols transmitted across the channel using some of the radio resource blocks [8].
 Afterward, the receiver uses the observed signals to determine a reliable CSI estimate.
-As the number of pilots scales with the number of users, the spectral efficiency decreases for higher number of users as less symbols are available for transmitting data.
+As the number of pilots scales with the number of users, the spectral efficiency decreases for a higher number of users as fewer symbols are available for transmitting data.
 To enhance channel estimation without increasing the number of pilot symbols, various methods have been developed that leverage the information embedded in the observed data symbols at the receiver to infer channel characteristics [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
@@
 The benefit of semi-blind channel estimation was first studied in [9], where Cramer-Rao bounds (CRBs) for blind, semi-blind, and training-based channel estimation were investigated in the context of single-input-multiple-output (SIMO) systems.
-In [10, 11], semi-blind channel estimation schemes based on maximum likelihood (ML) estimation were introduced.
-The asymptotic performance of the respective estimators was also studied in [10] for infinitely long data sequences.
+In [10, 11], semi-blind channel estimation schemes were introduced based on maximum likelihood (ML) estimation.
+The respective estimators’ asymptotic performances were also studied in [10] for infinitely long data sequences.
 Another asymptotic behavior, where the number of antennas grows to infinity, was studied in [12].
@@
 Early work on improving the least squares (LS) estimator using a semi-blind algorithm was conducted in [13].
-The authors in [14] propose to use partially decoded data for channel estimation.
 Furthermore, in [14], semi-blind and blind channel estimation was studied to enhance the maximum a-posteriori (MAP) channel estimates in massive MIMO systems.
 These findings are based on favorable propagation, which only holds for large antenna arrays deployed at the base station (BS).
-In [16], two semi-blind channel estimators based on the expectation maximization (EM) algorithm are studied.
+In [15], two semi-blind channel estimators are studied based on the expectation maximization (EM) algorithm.
 The assumption of a Gaussian distribution for the data was verified, leading to a closed-form solution for the E-step.
@@
 In [18], a data-aided iterative scheme is proposed for orthogonal time frequency space (OTFS) systems by employing affine-precoded superimposed pilots.
-The performance improvement achieved with these iterative approaches generally requires computational costly updates.
+The performance improvement achieved with these iterative approaches generally requires computationally costly updates.
 A low complexity iterative LS channel estimation algorithm is proposed in [19] for a massive MIMO turbo-receiver.
 In [20], the concept of semi-blind channel estimation was adapted for time domain synchronous-orthogonal frequency division multiplexing (OFDM) systems, where in addition to the pseudo noise sequence in the guard interval, the sent OFDM data symbols are exploited for the channel estimation.
-In [22], a framework was introduced for choosing reliable decoded data symbols, which can be interpreted as additional pilots.
-Similarly, reliably detected symbols are used for a semi-data-aided channel estimation in [23].
+In [21, 22, 23], different decision-directed frameworks were proposed, which treat reliably decoded data symbols as additional pilots.
+All of these methods utilize iterative approaches, thus exhibiting large complexities.
 In [24], peak-power carriers in an OFDM system are selected to eliminate the need to determine reliable data symbols at the receiver.
 Recently, a diffusion model-based approach for joint channel estimation and detection was proposed in [25], where a diffusion process is constructed that models the joint distribution of the channels and symbols given noisy observations.
-In this article, we focus on pilot-based estimators which minimize the MSE and investigate how these estimators can be extended for the semi-blind case.
-Notably, in [26] a subspace formulation of the MMSE estimator was used to mitigate pilot contamination in massive MIMO systems.
+Many state-of-the-art estimators fall in the class of unbiased estimators, which are potentially MSE-suboptimal.
+Instead, in this article, we focus on general, potentially biased, pilot-based estimators that minimize the MSE by utilizing prior information about the channel distribution, and investigate how these estimators can be extended for the semi-blind case.
+Contrarily to previous works, e.g., [21, 22, 23], we do not focus on using decoded data symbols as additional pilot but instead want to improve the estimators themselves.
+Although both approaches can be combined.
+Notably, in [26] a subspace formulation of the MMSE estimator was already used to mitigate pilot contamination in massive MIMO systems.
 The MMSE estimator is known to be the conditional mean estimator (CME) [27, Ch.
@@
 Since this true underlying distribution is unknown, machine learning methods rely on a representative data set, which is assumed to be available at the BS.
-Based on this data set the PDF of the user channels can be learned.
+Based on this data set, the user channels’ PDF can be learned.
 The first proposal of using a Gaussian mixture model (GMM) to formulate an estimator was done in [28] for the case of image processing.
 The approaches in [30, 31, 32] build on that by constructing a conditionally Gaussian latent model (CGLM) for the PDF of a BS cell environment.
-The learned CGLM not only enables MMSE channel estimation in [30, 31, 32] but can also be used for e.g., a limited feedback scheme as in [33].
+The learned CGLM not only enables MMSE channel estimation in [30, 31, 32] but can also be used for, e.g., a limited feedback scheme as in [33].
 In this work, we propose to utilize CGLMs to parameterize the CME in the semi-blind setting.
-The contributions of this work are summarized as follows:
-  • We introduce two variants of the linear MMSE (LMMSE) estimator incorporating subspace knowledge provided by the payload data symbols.
-First, we depict how the LMMSE channel estimator can solve a subspace estimation problem [26].
-As an alternative, we propose a projection method that is computationally more efficient since it allows for the pre-calculation of LMMSE filters.
-  • With theoretical derivations we show the superior MSE performance of the proposed projection method in the case of uncorrelated Rayleigh fading and perfect subspace knowledge.
+This work’s contributions are summarized as follows:
+  • We introduce two variants of the linear MMSE (MMSE) estimator incorporating subspace knowledge provided by the payload data symbols.
+Firstly, we depict how the LMMSE channel estimator can solve a subspace estimation problem [26].
+As an alternative, we propose a new projection method that is computationally more efficient since it allows for the pre-calculation of LMMSE filters.
+  • With theoretical derivations, we show the proposed projection method’s superior MSE performance in the case of uncorrelated Rayleigh fading and perfect subspace knowledge.
+  • We provide an asymptotical analysis of how the projected LMMSE outperforms the unbiased CRB for perfect subspaces.
+These derivations highlight the MMSE-based channel estimation’s superiority compared to unbiased channel estimation in the case of semi-blind channel estimation.
   • We show how the GMM [30] and variational autoencoder (VAE) [32], instances of the class of CGLMs, can be used to parameterize the semi-blind LMMSE estimator.
-  • Extensive simulations on different datasets, consisting of typical massive MIMO systems with multiple users and including real-world measurement data, show the superior performance of our proposed methods compared to state-of-the-art semi-blind channel estimation algorithms with respect to the MSE.
+  • Extensive simulations on different datasets, consisting of typical massive MIMO systems with multiple users and real-world measurement data, show our proposed methods’ superior performance compared to state-of-the-art semi-blind channel estimation algorithms in terms of the MSE.
 Preliminary results were presented in [1] and extended to the multi-user MIMO case in [34], which we extend further in the following aspects.
-The theoretical analyses in Section III enhance the foundation of the proposed semi-blind channel estimation strategies and provide analytic insights into the superior performance of the proposed projection method.
+The theoretical analyses in Section III enhance the proposed semi-blind channel estimation strategies’ foundation and provide analytic insights into the proposed projection method’s superior performance.
 We extend our concept of semi-blind MMSE channel estimation to the whole class of CGLMs, providing a more general framework to parameterize the semi-blind LMMSE estimator.
-Finally, we provide more comprehensive simulation results to show the strengths of our proposed strategies.
-[…]
+Finally, we provide more comprehensive simulation results to show our proposed strategies’ strengths.
+Notations: Matrices and vectors are denoted with boldface symbols.
+$\bm{0}$ and $\mathbf{I}_{N}$ denote the zero vector of appropriate size and the identity matrix of size $N\times N$, respectively.
+$\mathbb{E}[\cdot]$, $\mathrm{tr}(\cdot)$, and $\mathrm{range}(\cdot)$ denote the expectation, trace, and range operators, respectively.
+We use $(\cdot)^{\mathrm{T}}$, $(\cdot)^{\mathrm{H}}$, $(\cdot)^{-1}$, $(\cdot)^{\dagger}$ to denote the transpose, conjugate transpose, inverse, and pseudo inverse.
+$\|\cdot\|$ denotes the $2$-norm of a vector.
+$\mathcal{N}_{\mathbb{C}}(\bm{\mu},{\bm{C}})$ denotes the circularly symmetric complex Gaussian distribution with mean $\bm{\mu}$ and covariance matrix $\bm{C}$}.
 II System and Channel Model
-We consider a multi-user uplink system with $J$ single-antenna users and a BS equipped with $M$ receive antennas. The received signal vector at time instance $n$ is then
+We consider a multi-user system with $J$ single-antenna users and a BS equipped with $M$ receive antennas.
+We focus on the uplink scenario for this work, but the proposed methods can also be extended to the downlink.
+The received signal vector at time instance $n$ is then
 ${\bm{y}}(n)$ $={\bm{H}}{\bm{x}}(n)+{\bm{n}}(n),\quad n=1,...,N,$ (1)
-where ${\bm{x}}(n)=[x_{1}(n),...,x_{J}(n)]^{\mathrm{T}}\in\mathbb{C}^{J}$ and ${\bm{n}}(n)\in\mathbb{C}^{M}$ denote the signal sent by each of the $J$ users and the additive noise, respectively, whereas ${\bm{H}}=[{\bm{h}}_{1},...,{\bm{h}}_{J}]$ contains the individual channels of the users ${\bm{h}}_{j}\in\mathbb{C}^{M}$.
-The case of multiple antennas at the users can be transformed into (1) by viewing each active stream as a different user.
-The corresponding channel ${\bm{h}}_{j}$ can then be seen as the effective channel.
+where ${\bm{x}}(n)=[x_{1}(n),...,x_{J}(n)]^{\mathrm{T}}\in\mathbb{C}^{J}$ and ${\bm{n}}(n)\in\mathbb{C}^{M}$ denote the signal sent by each of the $J$ users and the additive noise, respectively, whereas ${\bm{H}}=[{\bm{h}}_{1},...,{\bm{h}}_{J}]$ contains the individual user channels ${\bm{h}}_{j}\in\mathbb{C}^{M}$.
+The case of multiple antennas at the users can be transformed into (1) by viewing each active stream as a different user with its corresponding channel ${\bm{h}}_{j}$ being the respective effective channel.
 For further details on semi-blind channel estimation in multi-user MIMO, we refer the reader to [34].
-For the task of channel estimation, we consider a channel coherence interval larger than the number of snapshots $N$, i.e., the channels are constant over all snapshots.
-We assume that the noise is Gaussian with ${\bm{n}}(n)\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M})$.
+For the case of downlink semi-blind channel estimation, a subspace of the effective channel needs to be identifiable to improve estimation performance.
+This is the case if not all degrees of freedom are utilized, e.g., fewer pilots than receive antennas.
+We consider a channel coherence interval larger than the number of snapshots $N$, i.e., the channels are constant over all snapshots.
+We assume that the noise is Gaussian with ${\bm{n}}(n)\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M})$ and the sent signals satisfy $\mathbb{E}[{\bm{x}}(n){\bm{x}}^{\mathrm{H}}(n)]=\frac{1}{J}\mathbf{I}_{J}$.
 In conventional channel estimation schemes, each user’s signals include $N_{p}$ uplink pilots. These pilots are known to the BS. Hence, the received observations at the BS side are
 ${\bm{Y}}=\left[{\bm{Y}}_{p}^{\prime},{\bm{Y}}_{d}\right]={\bm{H}}\left[{\bm{P}},{\bm{D}}\right]+{\bm{N}}={\bm{H}}{\bm{X}}+{\bm{N}},$ (2)
-where ${\bm{Y}}\in\mathbb{C}^{M\times N}$, ${\bm{Y}}^{\prime}_{p}\in\mathbb{C}^{M\times N_{p}}$, ${\bm{Y}}_{d}\in\mathbb{C}^{M\times N-N_{p}}$, ${\bm{P}}\in\mathbb{C}^{J\times N_{p}}$, and ${\bm{D}}\in\mathbb{C}^{J\times N-N_{p}}$ denote all received observations, received pilot observations, received payload data observations, sent pilots, and sent payload data symbols, respectively.
+where ${\bm{Y}}\in\mathbb{C}^{M\times N}$, ${\bm{Y}}^{\prime}_{p}\in\mathbb{C}^{M\times N_{p}}$, ${\bm{Y}}_{d}\in\mathbb{C}^{M\times N-N_{p}}$, ${\bm{P}}\in\mathbb{C}^{J\times N_{p}}$, and ${\bm{D}}=[{\bm{d}}_{1},\dots,{\bm{d}}_{J}]^{\mathrm{T}}$ denote all received observations, received pilot observations, received payload data observations, sent pilots, and sent payload data symbols, respectively.
+The $j$-th user’s payload symbols ${\bm{d}}_{j}\in\mathbb{S}_{j}^{N-N_{p}}$ include realizations of the used transmit constellation, where $\mathbb{S}_{j}$ denotes the set of constellation points of user $j$.
+In general, the employed user constellations can differ from each other.
 In order to fully illuminate the channels, the number of pilots is, at minimum, the number of users $N_{p}\geq J$, and orthogonal pilots are used.
-We set $N_{p}=J$, and utilize discrete Fourier transform (DFT) pilot sequences.
-After decorrelating the orthogonal pilot sequences the received pilot observations simplify to
-${\bm{Y}}_{p}={\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}={\bm{H}}{\bm{P}}{\bm{P}}^{\mathrm{H}}+{\bm{N}}{\bm{P}}^{\mathrm{H}}={\bm{H}}+{\bm{N}}_{p},$ (3)
+We set ${\bm{P}}{\bm{P}}^{\mathrm{H}}=\frac{N_{p}}{J}\mathbf{I}_{M}$, and utilize submatrices of the discrete Fourier transform (DFT) matrix of size $N_{p}\times N_{p}$.
+After decorrelating the orthogonal pilot sequences, the received pilot observations simplify to
+${\bm{Y}}_{p}={\bm{Y}}^{\prime}_{p}{\bm{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}={\bm{H}}{\bm{P}}{\bm{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}+{\bm{N}}{\bm{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}={\bm{H}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\frac{J}{N_{p}}}}{\bm{N}}_{p},$ (3)
 where ${\bm{N}}_{p}$ has the same statistics as $\bm{N}$} and, hence, we can omit the subscript.
-This enables to consider channel estimation from a per user perspective in the subsequent discussions.
+This decorrelation lets us consider channel estimation from a per-user perspective in the subsequent discussions.
 For reasons of simpler readability, the index for the respective user is, therefore, no longer given in the following.
-Consequently, we denote the pilot observation of a user as
+Consequently, we denote a user’s pilot observation as
 ${\bm{y}}_{p}={\bm{h}}+{\bm{n}},$ (4)
-with ${\bm{n}}\sim\mathcal{N}_{\mathbb{C}}(0,{\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M})$.
+with ${\bm{n}}\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{\bm{n}}=\sigma^{2}_{\mathrm{eff}}\mathbf{I}_{M})$, where $\sigma^{2}_{\mathrm{eff}}=\sigma^{2}\frac{J}{N_{p}}$ denotes the effective noise variance.
+Further, we assume normalized channels as $\mathbb{E}[\|{\bm{h}}\|^{2}]=M$, which lets us define the signal-to-noise ratio (SNR) as $\sigma^{-2}$.
+As the number of pilots $N_{p}$ results in a simple scaling of SNR in the pilot transmission phase, we use $N_{p}=J$ in the rest of this work.
 II-A Spatial Channel Model
@@
 based on a set of parameters $\bm{\delta}$}, which describe the directions and properties of the multi-path propagation clusters.
-The main angles are drawn independently from a uniform distribution in $[0,2\pi]$ and the path gains are independent zero-mean Gaussians.
+The central angles are drawn independently from a uniform distribution in $0,2\pi]$, and the path gains are independent zero-mean Gaussians.
 The spatial covariance matrix is given by
 ${\bm{C}}_{\bm{\delta}}=\int_{-\pi}^{\pi}g(\vartheta,{\bm{\delta}}){\bm{a}}(\vartheta){\bm{a}}^{\mathrm{H}}(\vartheta)\mathrm{d}\vartheta,$ (6)
-where $g(\vartheta,{\bm{\delta}})$ is the power density consisting of a weighted sum of Laplace densities, which have standard deviations $\sigma_{\text{AS}}$ corresponding to the angular spread of the propagation clusters.
-The BS employs a uniform linear array (ULA) with $M=64$ antennas and $\lambda/2$ spacing.
+where $g(\vartheta,{\bm{\delta}})$ is the power density consisting of a weighted sum of Laplace densities, which have standard deviations $\sigma_{\text{AS}}$ corresponding to the propagation clusters’ angular spread.
+The BS employs a uniform linear array (ULA) with $M$ antennas and $\lambda/2$ spacing.
 The steering vector is then given as
-${\bm{a}}(\vartheta)=\frac{1}{\sqrt{M}}\left[1,\mathrm{e}^{-\mathrm{j}\pi\sin(\vartheta)},\dots,\mathrm{e}^{-\mathrm{j}\pi(M-1)\sin(\vartheta)}\right]^{\mathrm{T}}.$ (7)
-For every channel sample, we consider a new ${\bm{\delta}}$ and draw the sample according to ${\bm{h}}\sim\mathcal{N}_{\mathbb{C}}\left(\bm{0},{\bm{C}}_{\bm{\delta}}\right)$.
+${\bm{a}}(\vartheta)=\frac{1}{\sqrt{M}}\left[1,\mathrm{e}^{-\mathrm{j}\pi\sin(\vartheta)},...,\mathrm{e}^{-\mathrm{j}\pi(M-1)\sin(\vartheta)}\right]^{\mathrm{T}}.$ (7)
+We consider a new $\bm{\delta}$} for every channel sample and draw the sample according to ${\bm{h}}\sim\mathcal{N}_{\mathbb{C}}\left(\bm{0},{\bm{C}}_{\bm{\delta}}\right)$.
 II-B Measurement Campaign
-Since synthetic data capture real-world CSI characteristics only up to some extent, we utilize real-world data from a measurement campaign conducted at the Nokia campus in Stuttgart, Germany, in October/November 2017, cf.
+Since synthetic data captures real-world CSI characteristics only up to some extent, we utilize real-world data from a measurement campaign conducted at the Nokia campus in Stuttgart, Germany, in October/November 2017, cf.
 [36].
 The BS antenna with a uniform rectangular array (URA) comprises $N_{v}=4$ vertical ($\lambda$ spacing) and $N_{h}=16$ horizontal ($\lambda/2$ spacing) single polarized patch antennas.
-The operating carrier frequency is $2.18$ GHz and the antenna was mounted on a rooftop approximately $20$ meters above the ground.
+The operating carrier frequency is $2.18$ GHz, and the antenna was mounted on a rooftop approximately $20$ meters above the ground.
 For further details, we refer the reader to [36].
@@
 In this section, we introduce two variants of the LMMSE estimator incorporating information provided by the payload data symbols.
-To do so, we first restrict ourselves to the case where perfect statistical knowledge is available at the receiver.
-In channel estimation, commonly, only the pilot observation ${\bm{y}}_{p}$ is considered for channel estimation. The MSE optimal estimator given the pilot observation ${\bm{y}}_{p}$ is the CME
+To do so, we restrict ourselves for now to the case where perfect statistical knowledge is available at the receiver.
+In channel estimation, commonly, only the pilot observation ${\bm{y}}_{p}$ is considered for channel estimation. The MSE-optimal estimator given the pilot observation ${\bm{y}}_{p}$ is the CME
 $\hat{{\bm{h}}}_{\text{CME}}=\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p}\right].$ (8)
@@
 where $\rho_{i}$ are the eigenvalues of $\bm{C}$}.
+In contrast to MMSE-based channel estimation, ML-based channel estimation does not utilize knowledge about the channels’ prior distribution, making it a suboptimal estimation problem.
 For our considerations concerning semi-blind channel estimation, we assume knowledge about the subspace defined by $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$, where we denote with $\bm{V}$} the left singular vectors of $\bm{H}$}, which span the same subspace as the columns of $\bm{H}$}.
-Generally, we can formulate the ML estimate of ${\bm{H}}$ in view of (2) as [16]
+Generally, we can formulate the (suboptimal) ML estimate of $\bm{H}$} in view of (2) as [15]
 $\min_{{\bm{H}},{\bm{X}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{h}}_{n}\right\|^{2}+\sum_{n=J+1}^{N}\left\|{\bm{y}}(n)-{\bm{H}}{\bm{x}}(n)\right\|^{2},$ (13)
-where the first term belongs to the pilot observations and the second part refers to the observation obtained from the payload symbols.
-In [16] the EM algorithm is introduced to solve this channel estimation problem in terms of maximum likelihood, whereas, in [11] a semi-blind method is derived based on utilizing the subspace $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$.
+where the first term belongs to the pilot observations, and the second part refers to the observation obtained from the payload symbols.
+Obtaining a closed-form solution for (13) is known to be hard [37].
+In [15], the EM algorithm is introduced to solve this channel estimation problem in terms of maximum likelihood, whereas, in [11], a semi-blind method is derived based on utilizing the subspace $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$.
 Now, given the subspace $\mathrm{range}({\bm{V}})$, we can reformulate (13) as [11]
-$\min_{{\bm{S}},{\bm{X}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{s}}_{n}\right\|^{2}+\sum_{n=J+1}^{N}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{S}}{\bm{x}}(n)\right\|^{2},$ (14)
-with ${\bm{H}}={\bm{V}}{\bm{S}}$ and ${\bm{S}}=[{\bm{s}}_{1},...,{\bm{s}}_{J}]\in\mathbb{C}^{J\times J}$. The optimal solution for ${\bm{X}}$ is given as
-${\bm{x}}^{*}(n)$ $=({\bm{S}}^{\mathrm{H}}{\bm{V}}^{\mathrm{H}}{\bm{V}}{\bm{S}})^{-1}{\bm{S}}^{\mathrm{H}}{\bm{V}}^{\mathrm{H}}{\bm{y}}(n)$ (15)
-$=({\bm{S}}^{\mathrm{H}}{\bm{S}})^{-1}{\bm{S}}^{\mathrm{H}}{\bm{V}}^{\mathrm{H}}{\bm{y}}(n).$ (16)
-Reinserting this solution into (13), the right part of the objective simplifies to
-$\min_{{\bm{S}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{s}}_{n}\right\|^{2}+\sum_{n=J+1}^{N}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}(n)\right\|^{2},$ (17)
-where only the left part depends on ${\bm{S}}$. Thus, the ML problem for the user of interest results in [11]
+$\min_{{\bm{S}},{\bm{X}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{s}}_{n}\right\|^{2}+\hskip-1.0pt\sum_{n=J+1}^{N}\hskip-1.0pt\left\|{\bm{y}}(n)-{\bm{V}}{\bm{S}}{\bm{x}}(n)\right\|^{2},$ (14)
+with ${\bm{H}}={\bm{V}}{\bm{S}}$ and ${\bm{S}}=[{\bm{s}}_{1},\dots,{\bm{s}}_{J}]\in\mathbb{C}^{J\times J}$. We can rewrite the second term in (14) as
+$\sum_{n=J+1}^{N}\hskip-5.0pt\left\|{\bm{y}}(n)-{\bm{V}}{\bm{S}}{\bm{x}}(n)\right\|^{2}$ $=\hskip-5.0pt\sum_{n=J+1}^{N}\hskip-5.0pt\left\|\left(\mathbf{I}_{M}-{\bm{V}}{\bm{V}}^{\mathrm{H}}\right){\bm{y}}(n)\right\|^{2}$
+$\quad+\left\|{\bm{V}}^{\mathrm{H}}{\bm{y}}(n)-{\bm{S}}{\bm{x}}(n)\right\|^{2},$ (15)
+where the equality comes from the fact that the two terms are orthogonal to each other.
+The first term in (15) is constant for given $\bm{V}$}.
+In the case of a continuous constellation for ${\bm{x}}(n)$, the second term vanishes by solving for $\bm{X}$}.
+Further, in the case of discrete symmetric signal constellations, the second term in (15) can not be jointly solved for $\bm{S}$} and $\bm{X}$} due to phase ambiguity.
+Thus, the ML problem for the user of interest reduces to the pilot observations as [11]
 $\min_{\bm{s}}\|{\bm{y}}_{p}-{\bm{V}}{\bm{s}}\|^{2},$ (16)
-with the closed form solution $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$. The MSE of this estimator is
+with the closed form solution $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$. This estimator’s MSE is
 $\mathrm{MSE}^{\mathrm{ML}}=\mathbb{E}\left[\|{\bm{h}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}\|^{2}\right]=J\sigma^{2}.$ (17)
@@
 Using the information in $\mathrm{range}({{\bm{V}}})$, we can solve the estimation within the subspace as previously proposed in [26].
-For this, the pilot system model in (4) is projected into the $J$-dimensional subspace as
+For this, the pilot system model in (4) is mapped into the $J$-dimensional subspace as
 ${\bm{y}}^{\prime}={\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$ $={\bm{V}}^{\mathrm{H}}{\bm{h}}+{\bm{V}}^{\mathrm{H}}{\bm{n}}={\bm{h}}^{\prime}+{\bm{n}}^{\prime}.$ (18)
@@
 $\hat{{\bm{h}}}^{\prime}=\;$ ${\bm{V}}^{\mathrm{H}}{\bm{C}}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}.$ (19)
-One should note that by design ${\bm{V}}$ actually depends on ${\bm{h}}$ and, hence, (21) is a suboptimal but feasible estimate for ${\bm{h}}^{\prime}$.
+By design, $\bm{V}$} depends on $\bm{h}$} and, hence, (19) is a suboptimal but feasible estimate for $\bm{h}$$^{\prime}$.
 After solving the estimation in the subspace for $\bm{h}$$^{\prime}$, the solution can be transformed back using [26]
@@
 III-B Projected Channel Estimator
-As an alternative approach, we propose the use of the orthogonal subspace projection ${\bm{P}}_{\bm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ as a preprocessing filter.
-Since the projector ${\bm{P}}_{\bm{H}}$ does not affect ${\bm{h}}$, the resulting projected observation is given by
+As an alternative approach, we propose using the orthogonal subspace projection ${\bm{P}}_{\bm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ as a preprocessing filter.
+Since the projector ${\bm{P}}_{\bm{H}}$ does not affect $\bm{h}$}, the resulting projected pilot observation is given by
 $\tilde{{\bm{y}}}={\bm{P}}_{\bm{H}}{\bm{y}}_{p}={\bm{h}}+{\bm{P}}_{\bm{H}}{\bm{n}}={\bm{h}}+\tilde{{\bm{n}}}$. (21)
-To formulate the CME $\hat{{\bm{h}}}=\mathbb{E}\left[{\bm{h}}\mid\tilde{{\bm{y}}}\right]$, we need to calculate the statistic of the noise $\tilde{{\bm{n}}}$ with
-${\bm{C}}_{\tilde{{\bm{n}}}}=\;$ $\mathbb{E}\left[\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right]=\mathbb{E}\left[\sigma^{2}{\bm{P}}_{\bm{H}}\right].$ (24)
-To get an intuitive understanding of (24), let us consider a scenario involving spatially uncorrelated channels, meaning that path gains and channel directions are uncorrelated.
-This is the case when users are uniformly distributed over the directions, e.g., the spatial channel model in Section II-A, resulting in a channel covariance matrix of the scenario that is a scaled identity [37, Def.
+To formulate the CME $\hat{{\bm{h}}}=\mathbb{E}\left[{\bm{h}}\mid\tilde{{\bm{y}}}\right$], we calculate the covariance matrix of $\tilde{{\bm{y}}}$ as
+$\mathbb{E}\left[\tilde{{\bm{y}}}\tilde{{\bm{y}}}^{\mathrm{H}}\right]$ $={\bm{C}}+\mathbb{E}\left[\sigma^{2}{\bm{P}}_{\bm{H}}\right],$ (22)
+where the mixing terms vanish due to the independence of $\bm{h}$} and $\bm{n}$}.
+To get an intuitive understanding of the noise covariance in (22), let us consider a scenario involving spatially uncorrelated channels, meaning that path gains and channel directions are uncorrelated.
+This is the case when users are uniformly distributed over the directions, e.g., the spatial channel model in Section II-A, resulting in a scenario’s channel covariance matrix that is a scaled identity [38, Def.
 2.3].
-In such a case, the matrices with the eigenvectors of the sample covariance matrix of such channels are distributed with Haar measure [38, Chap.
+In such a case, the matrices with the sample covariance matrix’s eigenvectors are distributed with Haar measure [39, Chap.
 1], i.e., uniformly distributed on the manifold of unitary matrices.
-Assuming spatially uncorrelated channels (24) results in
+Assuming spatially uncorrelated channels we have
 ${\bm{C}}_{\tilde{{\bm{n}}}}=\sigma^{2}\frac{J}{M}\mathbf{I}_{M},$ (23)
@@
 $\hat{{\bm{h}}}_{\text{proj}}={\bm{C}}\left({\bm{C}}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}\tilde{{\bm{y}}}$. (24)
-III-C Performance Analysis
+III-C Performance Analysis for Perfect Subspace Knowledge
 If (23) is true, the MSE of the proposed projected channel estimator can directly be written as (cf. Section -A)
@@
 $=\sum_{i=1}^{M}\frac{\rho_{i}\sigma^{2}}{\frac{M}{J}\rho_{i}+\sigma^{2}}.$ (26)
-Comparing the performance to the plain LMMSE we see that
+Comparing the performance to the plain LMMSE from (12) we see that
 $\frac{\rho_{i}\sigma^{2}}{\frac{M}{J}\rho_{i}+\sigma^{2}}\leq\frac{\rho_{i}\sigma^{2}}{\rho_{i}+\sigma^{2}},$ (27)
-holds for every $i=1,\dots,M$, resulting in $\mathrm{MSE}^{\mathrm{proj}}\leq\mathrm{MSE}^{\mathrm{plain}}$, cf.
-(12).
+holds for every $i\in\{1,\dots,M\}$, resulting in $\mathrm{MSE}^{\mathrm{proj}}\leq\mathrm{MSE}^{\mathrm{plain}}$.
 The inequality in (27) only holds with equality if $J=M$.
-Additionally, we can compare the MSE of the projected LMMSE to (19) by reformulating (28) as
+Additionally, we can compare the MSE of the projected LMMSE to the ML estimator from (17) by reformulating (26) as
 $\mathrm{MSE}^{\mathrm{proj}}=\frac{J}{M}\sigma^{2}\sum_{i=1}^{M}\frac{\rho_{i}}{\rho_{i}+\frac{J}{M}\sigma^{2}}\leq J\sigma^{2}=\mathrm{MSE}^{\text{ML}}.$ (28)
-To compare the projected variant to the subspace LMMSE, let us consider uncorrelated Rayleigh fading such that the channel covariance matrix is given as ${\bm{C}}=\mathbf{I}_{M}$.
-The MSE of the projected variant results in
+In the case of uncorrelated Rayleigh fading, the channel covariance matrix is given as ${\bm{C}}=\mathbf{I}_{M}$. The projected variant’s MSE results in
 $\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}=\frac{JM\sigma^{2}}{{M}+J\sigma^{2}}.$ (29)
-For the case of ${\bm{C}}=\mathbf{I}_{M}$, the subspace LMMSE estimator boils down to
+Now let us compare the projected variant to the subspace LMMSE. For the case of ${\bm{C}}=\mathbf{I}_{M}$, the subspace LMMSE estimator boils down to
 $\hat{{\bm{h}}}_{\mathrm{sub}}=\frac{1}{1+\sigma^{2}}{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}},$ (30)
@@
 Fig.
-1 shows the performance of the individual channel estimators based on perfect genie knowledge for the case of uncorrelated Rayleigh fading.
-As one can see, the projected channel estimator outperforms all other estimators across the whole signal-to-noise ratio (SNR) range.
+1a shows the individual channel estimators’ performances based on perfect subspace knowledge for the case of uncorrelated Rayleigh fading.
+As one can see, the projected channel estimator outperforms all other estimators across the whole SNR range.
 Further, we realize that the subspace LMMSE converges to the ML method from above.
-Figure 1: MSE over the SNR for given channel estimations based on perfect subspace and perfect statistical knowledge in a $J=8$ user and $M=64$ antennas scenario with uncorrelated Rayleigh fading.
+Additionally, performance guarantees of the proposed projected LMMSE compared to the unbiased CRB as derived in [15] can be given.
+We formulate the following theorem for the asymptotic region, where $M$ and $N$ tend to infinity.
+Theorem 1.
+We consider the limits of $N\to\infty$ and $M/N\to\alpha$ with $\alpha>0$, while keeping the number of pilot symbols in the order of users $J$.
+Then, the projected LMMSE based on perfect knowledge of the subspace spanned by the left singular vectors of the channel matrix $\bm{H}$} is equal to or better than any unbiased estimator in terms of MSE for uncorrelated Rayleigh fading.
+Proof.
+We take the limit of the MSE of the projected LMMSE estimator as
+$\lim_{\begin{subarray}{c}N\to\infty\\ \frac{M}{N}\to\alpha\end{subarray}}\mathrm{MSE}^{\mathrm{proj}}$ $=\lim_{\begin{subarray}{c}N\to\infty\\ \frac{M}{N}\to\alpha\end{subarray}}\frac{J}{M}\sigma^{2}\sum_{i=1}^{M}\frac{\rho_{i}}{\rho_{i}+\frac{J}{M}\sigma^{2}}$ (32)
+$=J\sigma^{2}\leq(1+\alpha)J\sigma^{2}=\mathrm{CRB}_{\mathrm{iid,d}},$ (33)
+where $\mathrm{CRB}_{\mathrm{iid,d}}$ denotes the deterministic CRB’s limit for uncorrelated Rayleigh fading.
+The last equality in (33) was proven in [40, Theorem 4].
+Furthermore, it is stated in [40] that in this asymptotic limit, the deterministic CRB is always lower than the stochastic CRB.
+∎
+Theorem 1 provides some insights on the asymptotic superiority of our proposed projected LMMSE strategy based on perfect subspace knowledge compared to any unbiased channel estimator.
+Additionally, the MSE of the subspace variant as well as the plain LMMSE approach infinity for the considered limit, making them irrelevant for this asymptotic regime.
+Interestingly, the projected LMMSE also approaches $J\sigma^{2}$ for the case where only $M$ tends to infinity, making the projected LMMSE a valid candidate for large antenna systems as the performance no longer depends on $M$.
+III-D Maximum Likelihood Subspace Estimation
+After introducing the methods utilizing the additional subspace information provided by $\mathrm{range}({\bm{V}})$ to enhance the CSI estimation quality, let us consider the estimation of such a subspace.
+As the received data symbols are transmitted over the same channels, we can use these payload symbols to estimate the subspace containing all user channels.
+To this end, let us reconsider the ML estimate of $\bm{H}$} in (13).
+Instead of directly optimizing on this ML formulation as done in [15], which generally does not result in the MMSE, we only take this log-likelihood formulation as an intermediate step to estimate the subspace $\mathrm{range}({\bm{V}})$.
+First, let us consider the payload ML objective formulation in (15).
+We can then reformulate the problem, neglecting the second term due to phase ambiguity, resulting in [11]
+$\max_{\bm{H}}\;\mathrm{tr}\left({\bm{P}}_{\bm{H}}\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}\right)$, (34)
+where ${\bm{P}}_{\bm{H}}={\bm{H}}({\bm{H}}^{\mathrm{H}}{\bm{H}})^{-1}{\bm{H}}^{\mathrm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ and $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N-J}{\bm{Y}}_{d}{\bm{Y}}_{d}^{\mathrm{H}}$, with ${\bm{Y}}_{d}$ from (2).
+The maximization in (34) is solved by setting ${\bm{P}}_{\bm{H}}$ equal to $\hat{{\bm{V}}}_{d}\hat{{\bm{V}}}_{d}^{\mathrm{H}}$ with $\hat{{\bm{V}}}_{d}$ holding the $J$ dominant eigenvectors of the receive sample covariance matrix $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}$.
+Additionally, it is trivial to see that the first term in (13) is minimized by ${\bm{h}}_{n}={\bm{y}}(n)$.
+The subspace spanned by the solution ${\bm{h}}_{n}={\bm{y}}(n)$ is the same as the subspace spanned by the $J$ eigenvectors of the sample covariance matrix $\hat{{\bm{C}}}^{(p)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{J}{\bm{Y}}_{p}{\bm{Y}}_{p}^{\mathrm{H}}$, which ignores the additional phase information contained in the pilot observation.
+Thus, the overall subspace estimate $\hat{{\bm{V}}}=[{\bm{v}}_{1},\dots,{\bm{v}}_{J}]$ is found by taking the $J$ dominant eigenvectors ${\bm{v}}_{j}$ of the sample covariance matrix defined as
+$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}.$ (35)
+This result has also been used in [26].
+To save computational complexity while calculating the sample covariance matrix in (35), information from the previous coherence intervals can be utilized.
+This can be done by adaptively updating the subspace using efficient tracking algorithms as proposed in, e.g., [41, 42].
+(a) Perfect Subspace Knowledge
+(b) Subspace Estimation based on $N=200$ Snapshots
+Figure 1: MSE over the SNR for given channel estimations based on perfect statistical knowledge in a $J=8$ user and $M=64$ antennas scenario with uncorrelated Rayleigh fading and Gaussian symbols using (a) perfect subspace knowledge and (b) $N=200$ snapshots for subspace estimation.
+III-E Performance Analysis for Estimated Subspace
+Fig.
+1b shows the individual channel estimators’ performances based on an estimated subspace using $N=200$ snapshots for uncorrelated Rayleigh fading.
+Additionally, the unbiased CRBs derived in [15] and their extensions to the Bayesian Cramer-Rao bounds (BCRBs), derived in Section -C, are shown.
+For increasing SNR, all semi-blind channel estimators considered in this section converge to the CRBs and BCRBs.
+Furthermore, we see that for low SNR values the introduced MMSE-based semi-blind channel estimators outperform the different unbiased CRBs, which are lower bounds for, e.g., the EM-based estimator from [15].
+In order to quantify the error introduced by the subspace estimation in comparison to perfect subspace knowledge, let us consider the projected LMMSE as introduced in Section III-B in combination with the subspace estimation from Section III-D.
+We assume Gaussian distributed symbols for the following derivations.
+The introduced error by estimating the subspace can be bounded using Theorem 2.
+Theorem 2.
+Let the subspace spanned by the left singular vectors of the channel matrix $\bm{H}$} be estimated based on $N$ observations.
+The probability that the difference between the projected LMMSE channel estimators utilizing the estimated and true subspace, respectively, is less than $\varepsilon$ is given as
+$P\left(\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]\leq\varepsilon\right)\\ \geq 1-\frac{4k^{2}(M-J)\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})}{J^{2}N\varepsilon}\sum_{j\leq J}\frac{|s_{\max}|^{2}+2\sigma^{2}}{\lambda_{j}^{2}},$ (36)
+where $\lambda_{j}$ is the $j$-th eigenvalue of ${\bm{H}}{\bm{H}}^{\mathrm{H}}$, $s_{\max}=\max_{j\in\{1,\dots,J\}}s_{j}$ with ${\bm{h}}={\bm{V}}{\bm{s}}$, and
+$k^{2}=\sigma^{2}\left(\sigma^{2}+\mathrm{tr}({\bm{C}}_{{\bm{y}}\mid{\bm{H}}})\right).$ (37)
+Proof.
+See Appendix -D
+The probability bound in Theorem 2 is very loose, as the used inequalities are not tight.
+Nevertheless, this bound can be used to show the superiority of the projected LMMSE based on an estimated subspace for a high number of observations compared to the other considered estimators.
+Table I gives the bound values for the case of $N=10^{4}$ observations at SNR =0=0 dB for different number of users.
+As one can see, the projected LMMSE outperforms the other estimators with high probability.
+TABLE I: Lower bound on the probability of the projected LMMSE outperforming the other estimators for $N=10^{4}$ observations at SNR =0=0dB.
+Number of users $J$ 88 $16$ $32$
+$P(\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}\leq\mathrm{MSE}^{\mathrm{sub}}_{\mathrm{iid}})$ $\geq 0.9873$ $\geq 0.9791$ $\geq 0.9642$
+$P(\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}\leq\mathrm{MSE}^{\mathrm{ML}}_{\mathrm{iid}})$ $\geq 0.8453$ $\geq 0.9539$ $\geq 0.9907$
+$P(\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}\leq\mathrm{MSE}^{\mathrm{plain}}_{\mathrm{iid}})$ $\geq 0.9946$ $\geq 0.9923$ $\geq 0.9907$
 IV Proposed Semi-Blind Channel Estimation - Utilizing Generative Prior
-In practice, (10) can not be utilized directly as the channels have to be Gaussian distributed, and this distribution needs to be known.
-In general, the CME given the pilot observation ${\bm{y}}_{p}$ is formulated as
+In practice, the user channels do not follow a Gaussian distribution. Hence, (10) can not be utilized directly and the CME given the pilot observation ${\bm{y}}_{p}$ is formulated as
 $\hat{{\bm{h}}}_{\text{CME}}=\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p}\right]=\int{\bm{h}}\frac{p_{{\bm{n}}}({\bm{y}}_{p}-{\bm{h}})p({\bm{h}})}{p({\bm{y}}_{p})}\mathrm{d}{\bm{h}}.$ (38)
-As can be seen in (34), the CME generally can not be computed analytically.
+As seen in (38), the CME generally can not be computed analytically.
 First, the CME needs access to $p({\bm{h}})$, which is generally unavailable in practice.
@@
 $=\int\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p},{\bm{c}}\right]p({\bm{c}}\mid{\bm{y}}_{p})\mathrm{d}{\bm{c}}$ (41)
-$\approx\int\hat{{\bm{h}}}_{\bm{c}}({\bm{y}}_{p})p({\bm{c}}\mid{\bm{y}}_{p})\mathrm{d}{\bm{c}}$ (38)
+$=\int\hat{{\bm{h}}}_{\bm{c}}({\bm{y}}_{p})p({\bm{c}}\mid{\bm{y}}_{p})\mathrm{d}{\bm{c}}$ (42)
 where
@@
 denotes the LMMSE estimate given $\bm{c}$}.
-However, finding a suitable condition $\bm{\delta}$ can be challenging, in particular as the true distribution of ${\bm{h}}$ is unknown.
+However, finding a suitable condition $\bm{c}$} can be challenging, particularly as the true distribution of $\bm{h}$} is unknown.
 To this end, CGLMs were proposed in [30, 31, 32], which approximate the CME based on a GMM, mixtures of factor analyzers (MFA), and VAE, respectively.
 All three methods learn a model that provides the conditional Gaussian distribution ${\bm{h}}\mid{\bm{c}}$ based on a discrete (GMM, MFA) or continuous (VAE) latent variable $\bm{c}$}.
-In this work, we focus on the GMM and VAE, which we adapt to semi-blind channel estimation in the following.
+This work focuses on the GMM and VAE, which we adapt to semi-blind channel estimation in the following.
 IV-A GMM-based Semi-blind Channel Estimation
+Algorithm 1 Subspace GMM Channel Estimator
+Offline Training Phase
+1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
+2:Fit the GMM with the EM algorithm, cf. [30]
+3:
+[…]
+[…]
+[…]
+[…]
+8:for $j=1,\dots,J$ do
+9:  ${\bm{y}}^{\prime}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
+10:  for $k=1,\dots,K$ do
+11:   $\hat{{\bm{h}}}^{\prime}_{k}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{y}}^{\prime}$
+12:  end for
+13:  $\hat{{\bm{h}}}_{j}\leftarrow{\bm{V}}\sum_{k=1}^{K}p(k\mid{\bm{y}}^{\prime})\hat{{\bm{h}}}^{\prime}_{k}$
+14:end for
+15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
+Algorithm 2 Projected GMM Channel Estimator
+Offline Training Phase
+1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
+2:Fit the GMM with the EM algorithm, cf. [30]
+3:
+[…]
+[…]
+[…]
+[…]
+8:for $j=1,\dots,J$ do
+9:  $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
+10:  for $k=1,\dots,K$ do
+11:   $\hat{{\bm{h}}}_{k}\leftarrow{\bm{C}}_{k}\left({\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}\tilde{{\bm{y}}}$
+12:  end for
+13:  $\hat{{\bm{h}}}_{j}\leftarrow\sum_{k=1}^{K}p(k\mid\tilde{{\bm{y}}})\hat{{\bm{h}}}_{k}$
+14:end for
+15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
 Based on the universal approximation property of GMMs [44], the PDF of $\bm{h}$} is approximated by
 $f_{\bm{h}}^{(K)}({\bm{h}})=\sum_{k=1}^{K}p(k)\mathcal{N}_{\mathbb{C}}({\bm{h}};\bm{\mu}_{k},{\bm{C}}_{k}),$ (44)
-where $p(k)$, $\bm{\mu}_{k}$ and ${\bm{C}}_{k}$ are the mixing coefficients, means and covariance matrices of the $k$-th GMM component, respectively.
-As we are considering wireless channels, the mean of each component is set to $\bm{\mu}_{k}=\bm{0}$, cf.
+where $p(k)$, $\bm{\mu}_{k}$, and ${\bm{C}}_{k}$ are the mixing coefficients, means, and covariance matrices of the $k$-th GMM component, respectively.
+As we consider wireless channels, each component’s mean is set to $\bm{\mu}_{k}=\bm{0}$, cf.
 [43].
-The fitting of the components in (40) is accomplished with the well-known EM algorithm [41] based on a set =$\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ of $T$ channel samples as training data.
-Based on the formulation in (40) the conditional PDF given $k$ is
+The fitting of the components in (44) is accomplished with the well-known EM algorithm [45] based on a set $\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ of $T$ channel samples as training data.
+Based on the formulation in (44) the conditional PDF is
 ${\bm{h}}\mid k\sim\mathcal{N}_{\mathbb{C}}({\bm{h}};\bm{0},{\bm{C}}_{k}).$ (45)
-Thus, in the case of a GMM we have a discrete latent variable, which helps us parameterize the CME. The resulting semi-blind subspace GMM can be formulated as
+Thus, we have a discrete latent variable in the case of a GMM, which helps us parameterize the CME. The resulting semi-blind subspace GMM can be formulated as
 $\hat{{\bm{h}}}_{\text{sub. GMM}}={\bm{V}}\hat{{\bm{h}}}^{\prime}_{\text{GMM}}={\bm{V}}\sum_{k=1}^{K}p(k\mid{\bm{y}}^{\prime})\hat{{\bm{h}}}^{\prime}_{\text{GMM},k},$ (46)
@@
 and the associated responsibilities
-$p(k\mid\tilde{{\bm{y}}})=\frac{p(k)\mathcal{N}_{\mathbb{C}}\left(\tilde{{\bm{y}}};\bm{0},{\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)}{\sum_{i=1}^{K}p(i)\mathcal{N}_{\mathbb{C}}\left(\tilde{{\bm{y}}};\bm{0},{\bm{C}}_{i}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)}.$ (47)
+$p(k\mid\tilde{{\bm{y}}})=\frac{p(k)\mathcal{N}_{\mathbb{C}}\left(\tilde{{\bm{y}}};\bm{0},{\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)}{\sum_{i=1}^{K}p(i)\mathcal{N}_{\mathbb{C}}\left(\tilde{{\bm{y}}};\bm{0},{\bm{C}}_{i}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)}$.
+(51)
 The respective estimators are summarized in Algorithm 1 and Algorithm 2.
-Algorithm 1 Subspace GMM Channel Estimator
-Offline Training Phase
-1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
-2:Fit the GMM with the EM algorithm, cf. [30]
-3:
-4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
-5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
-6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
-7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
-8:for $j=1,\dots,J$ do
-9:     ${\bm{y}}^{\prime}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
-10:     for $k=1,\dots,K$ do
-11:         $\hat{{\bm{h}}}^{\prime}_{k}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{y}}^{\prime}$
-12:     end for
-13:     $\hat{{\bm{h}}}_{j}\leftarrow{\bm{V}}\sum_{k=1}^{K}p(k\mid{\bm{y}}^{\prime})\hat{{\bm{h}}}^{\prime}_{k}$
-14:end for
-15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
-Algorithm 2 Projected GMM Channel Estimator
-Offline Training Phase
-1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
-2:Fit the GMM with the EM algorithm, cf. [30]
-3:
-4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
-5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
-6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
-7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
-8:for $j=1,\dots,J$ do
-9:     $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
-10:     for $k=1,\dots,K$ do
-11:         $\hat{{\bm{h}}}_{k}\leftarrow{\bm{C}}_{k}\left({\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}\tilde{{\bm{y}}}$
-12:     end for
-13:     $\hat{{\bm{h}}}_{j}\leftarrow\sum_{k=1}^{K}p(k\mid\tilde{{\bm{y}}})\hat{{\bm{h}}}_{k}$
-14:end for
-15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
+IV-B VAE-based Semi-blind Channel Estimation
 Figure 2: Structure of a semi-blind VAE. The matrix $\hat{{\bm{V}}}$ contains the $J$ dominant eigenvectors of (35). The encoder and decoder represent DNNs.
-IV-B VAE-based Semi-blind Channel Estimation
-To learn the unknown distribution $f_{\bm{h}}({\bm{h}})$ using a VAE, we lower bound the parameterized likelihood $p_{\bm{\theta}}({\bm{h}})$ using the evidence-lower bound (ELBO).
-To formulate the ELBO, the variational distributions $q_{\bm{\phi}}({\bm{z}}\mid{\bm{y}}^{\prime})$ and $q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$ are introduced, which approximate $p({\bm{z}}\mid{\bm{y}}^{\prime})$ and $p({\bm{z}}\mid\tilde{{\bm{y}}})$, respectively.
-In contrast to the GMM, the used subspace $\mathrm{range}({\bm{V}})$ is unknown to the encoder of the VAE making $p({\bm{z}}\mid{\bm{y}}^{\prime})$ difficult to learn.
-Additionally, the dimension of the encoder input would depend on the number of users in the system.
-Thus, we propose to approximate both posteriors $p({\bm{z}}\mid{\bm{y}}^{\prime})$ and $p({\bm{z}}\mid\tilde{{\bm{y}}})$ with $q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$.
-A version of the ELBO for this case, which is accessible, can be written as [42]
+[…]
 $\mathcal{L}_{{\bm{\theta}},{\bm{\phi}}}=\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}})]-\mathrm{D}_{\mathrm{KL}}(q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})\mid\mid p({\bm{z}})),$ (52)
-where $\mathbb{E}_{q_{\bm{\phi}}}[\cdot]=\mathbb{E}_{q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})}[\cdot]$ is the expectation over the variational distribution $q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$.
-The second term in (48) is the Kullback-Leibler (KL) divergence
+[…]
 $\mathrm{D}_{\mathrm{KL}}(q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})\mid\mid p({\bm{z}}))=\mathbb{E}_{q_{\bm{\phi}}}\left[\log\left(\frac{q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})}{p({\bm{z}})}\right)\right].$ (53)
@@
 $p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}})$ $=\mathcal{N}_{\mathbb{C}}(\bm{\mu}_{\bm{\theta}}({\bm{z}}),{{\bm{C}}}_{\bm{\theta}}({\bm{z}})),$ (54)
-$q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$ $=\mathcal{N}(\bm{\mu}_{\bm{\phi}}(\tilde{{\bm{y}}}),\mathrm{diag}(\bm{\sigma}^{2}_{\bm{\phi}}(\tilde{{\bm{y}}}))).$
+[…]
 The resulting semi-blind VAE structure is shown in Fig.
 2.
-In the case of a ULA or URA at the BS the channel covariance matrix is either Toeplitz or block-Toeplitz, respectively.
-As shown in [39], the conditional covariance matrix at the output of the VAE preserves this structure.
+In the case of a ULA or URA at the BS, the channel covariance matrix is either Toeplitz or block-Toeplitz, respectively.
+As shown in [43], the conditional covariance matrix at the VAE’s output preserves this structure.
 Thus, we parameterize the output covariance matrix as
@@
 [47].
-Further, for the (block-)Toeplitz parameterization we can set $\bm{\mu}_{\bm{\theta}}({\bm{z}})=\bm{0}$, cf.
+Further, for the (block-)Toeplitz parameterization, we can set $\bm{\mu}_{\bm{\theta}}({\bm{z}})=\bm{0}$, cf.
 [43].
-After successfully training the VAE, the output is a local parameterization of $f_{\bm{h}}({\bm{h}})$ as conditionally Gaussian
-${\bm{h}}\mid{\bm{z}}\sim p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}}).$ (52)
-As analyzed in [32] it is a reasonable approximation to set
-$p({\bm{z}}\mid\tilde{{\bm{y}}})=\begin{cases}1\quad\text{if }{\bm{z}}=\bm{\mu}_{\bm{\phi}}(\tilde{{\bm{y}}}),\\ 0\quad\text{otherwise}.\end{cases}$ (53)
-Based on this parameterization we can formulate the semi-blind VAE-based estimators as
-[…]
-and
-[…]
-×$\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}).$ (55)
-The respective estimators are summarized in Algorithm 3 and Algorithm 4.
-For a more detailed introduction into the VAE framework and its usage for parameterization of the CME we refer the reader to [32].
 Algorithm 3 Subspace VAE Channel Estimator
 Offline Training Phase
-1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
+1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
 2:Fit the VAE by optimizing the ELBO, cf. [32]
 3:
-4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
-5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
-6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
-7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
+[…]
+[…]
+[…]
+[…]
 8:for $j=1,\dots,J$ do
 9:  $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
-10:     $\bm{\mu}_{\bm{\theta}}({\bm{z}}),{\bm{C}}_{\bm{\theta}}({\bm{z}})\leftarrow\mathrm{VAE}(\tilde{{\bm{y}}})$
+[…]
 11:  $\hat{{\bm{h}}}_{j}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}$
-12:           ×$\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}})$
+12:     $\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}})$
 13:end for
@@
 Offline Training Phase
-1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
+1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
 2:Fit the VAE by optimizing the ELBO, cf. [32]
 3:
-4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
-5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
-6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
-7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
+[…]
+[…]
+[…]
+[…]
 8:for $j=1,\dots,J$ do
 9:  $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
-10:     $\bm{\mu}_{\bm{\theta}}({\bm{z}}),{\bm{C}}_{\bm{\theta}}({\bm{z}})\leftarrow\mathrm{VAE}(\tilde{{\bm{y}}})$
+[…]
 11:  $\hat{{\bm{h}}}_{j}\leftarrow\bm{\mu}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\bm{\theta}}({\bm{z}})\left({\bm{C}}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}}))$
-12:           ×$\times(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}}))$
+12:     $\times(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}}))$
 13:end for
 14:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
-IV-C Maximum Likelihood Subspace Estimation
-After introducing the methods utilizing the additional subspace information provided by $\mathrm{range}({\bm{V}})$ to enhance the CSI estimation quality, let us consider the estimation of such a subspace.
-As the received data symbols are also transmitted over the same channel, they can be used to estimate the subspace containing all user channels.
-To this end, let us reconsider the ML estimate of ${\bm{H}}$ in (13).
-Instead of directly optimizing on this ML formulation as done in [16], which generally does not result in the MMSE, we only take this log-likelihood formulation as an intermediate step to estimate the subspace $\mathrm{range}({\bm{V}})$.
-First, let us consider the right term of the objective function in (13).
-We can then reformulate the problem by again solving for ${\bm{X}}$ first and reinserting the solution resulting in [11]
-$\max_{\bm{H}}\;\mathrm{tr}\left({\bm{P}}_{\bm{H}}\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}\right),$ (56)
-where ${\bm{P}}_{\bm{H}}={\bm{H}}({\bm{H}}^{\mathrm{H}}{\bm{H}})^{-1}{\bm{H}}^{\mathrm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ and $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N-J}{\bm{Y}}_{d}{\bm{Y}}_{d}^{\mathrm{H}}$, with ${\bm{Y}}_{d}$ from (2).
-The maximization in (56) is solved by setting ${\bm{P}}_{\bm{H}}$ equal to $\hat{{\bm{V}}}_{d}\hat{{\bm{V}}}_{d}^{\mathrm{H}}$ with $\hat{{\bm{V}}}_{d}$ holding the $J$ dominant eigenvectors of the receive sample covariance matrix $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}$.
-This result has also been used in [26].
-Additionally, it is trivial to see that the first term in (13) is minimized by ${\bm{h}}_{n}={\bm{y}}(n)$.
-The subspace spanned by the solution ${\bm{h}}_{n}={\bm{y}}(n)$ is the same as the subspace spanned by the $J$ eigenvectors of the sample covariance matrix $\hat{{\bm{C}}}^{(p)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{J}{\bm{Y}}_{p}{\bm{Y}}_{p}^{\mathrm{H}}$, which ignores the additional phase information contained in the pilot observation.
-Thus, the overall subspace estimate $\hat{{\bm{V}}}=[{\bm{v}}_{1},\dots,{\bm{v}}_{J}]$ is found by taking the $J$ dominant eigenvectors ${\bm{v}}_{j}$ of the sample covariance matrix defined as
-$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}.$ (57)
-To utilize information from the previous coherence intervals, one can adaptively update the subspace using efficient tracking algorithms as proposed in, e.g., [44, 45].
-IV-D Complexity Analysis
+After successfully training the VAE, the output is a local parameterization of $p({\bm{h}})$ as conditionally Gaussian
+${\bm{h}}\mid{\bm{z}}\sim p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}}).$ (56)
+As analyzed in [32], it is a reasonable approximation to set
+$p({\bm{z}}\mid\tilde{{\bm{y}}})=\begin{cases}1\quad\text{if }{\bm{z}}=\bm{\mu}_{\bm{\phi}}(\tilde{{\bm{y}}}),\\ 0\quad\text{otherwise}.\end{cases}$ (57)
+Based on this parameterization, we can formulate the semi-blind VAE-based estimators as
+$\hat{{\bm{h}}}_{\text{proj.
+VAE}}=\;$ $\bm{\mu}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\bm{\theta}}({\bm{z}})\left({\bm{C}}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}})),$ (58)
+and
+$\hat{{\bm{h}}}_{\text{sub.
+VAE}}=\;$ ${\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}$
+$\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}).$ (59)
+The respective estimators are summarized in Algorithm 3 and Algorithm 4. For a more detailed introduction into the VAE framework and its usage for CME parameterization, we refer the reader to [32].
+(a) QPSK
+(b) 16QAM
+(c) Gaussian Symbols
+Figure 3: NMSE over the SNR for given channel estimations in a $J=8$ user and $M=64$ antennas scenario based on $N=200$ observations, including $N_{p}=J$ pilots (if not stated otherwise) and (a) QPSK, (b) 16QAM, and (c) Gaussian data symbols, using measurement data (Sec.
+II-B).
+IV-C Complexity Analysis
 The standalone GMM estimator proposed by [30] precomputes the filters used for the individual components, resulting in a complexity of $\mathcal{O}(KM^{2})$.
-For the standalone VAE the complexity is given as $\mathcal{O}(DM^{2})$ [32], where $D$ denotes the number of layers in the forward pass of the VAE.
-For our semi-blind methods, the calculation of the subspace requires $\mathcal{O}((N+J)M^{2})$.
+For the standalone VAE, the complexity is given as $\mathcal{O}(DM^{2})$ [32], where $D$ denotes the number of layers in the VAE’s forward pass.
+For our semi-blind methods, the subspace calculation requires $\mathcal{O}((N+J)M^{2})$.
 This results from calculating the sample covariance matrix with $\mathcal{O}(NM^{2})$ and taking the eigenvectors of the $J$ largest eigenvalues for the solution of (34).
-Using the projection approximation subspace tracking (PAST) algorithm [45], the computational complexity of calculating the subspace reduces to $\mathcal{O}(JM)$ for every update.
-In the case of the subspace GMM the $K$ LMMSE estimates can not be precomputed, which results in a complexity of $\mathcal{O}(K(M^{2}+JM^{2}+J^{3}))$.
+Using the projection approximation subspace tracking (PAST) algorithm [42], the computational complexity of the subspace calculation reduces to $\mathcal{O}(JM)$ for every update.
+In the case of the subspace GMM, the $K$ LMMSE estimates can not be precomputed, which results in a complexity of $\mathcal{O}(K(M^{2}+JM^{2}+J^{3}))$.
 Similarily, the subspace VAE exhibits a complexity of $\mathcal{O}(DM^{2}+JM^{2}+J^{3})$.
@@
 One should note that the calculation for each of the $K$ components in the GMM can be parallelized.
-Similarly, the computations in the convolutional layers of the VAE can be parallelized, mitigating the complexity.
+Similarly, the computations in the VAE’s convolutional layers can be parallelized, mitigating the complexity.
 V Baseline estimators
-To compare our methods, the following baseline channel estimators are considered.
+The following baseline channel estimators are considered for comparison with our methods.
 Based on the found subspace $\mathrm{range}({\bm{V}})$ we can formulate the pilot-based ML estimator as $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$, which is the closed-form solution to (16).
 This can be interpreted as the subspace-adjusted version of the conventional least squares (LS) channel estimator given as $\hat{{\bm{h}}}_{\text{LS}}={\bm{y}}_{p}$.
-Another estimator is based on the sample covariance matrix, which we can compute from the training data set ${\mathcal{H}}$ to infer the global statistics of the channels as
+Another estimator is based on the sample covariance matrix, which we can compute from the training data set $\mathcal{H}$} to infer the channels’ global statistics as
 ${\bm{C}}_{s}=\frac{1}{|{\mathcal{H}}|}\sum_{{{\bm{h}}}\in{\mathcal{H}}}{{\bm{h}}}{{\bm{h}}}^{\mathrm{H}}.$ (60)
-We can use the matrix ${\bm{C}}$ as statistical prior to parameterize the semi-blind channel estimators outlined in Section III-A and Section III-B as
+We can use the matrix ${\bm{C}}_{s}$ as a statistical prior to parameterize the semi-blind channel estimators outlined in Section III-A and Section III-B as
 $\hat{{\bm{h}}}_{\text{sub.
@@
 $\hat{{\bm{h}}}_{\text{proj. s-cov}}={\bm{C}}_{s}\left({\bm{C}}_{s}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}{\bm{P}}_{\bm{H}}{\bm{y}}_{p}.$ (62)
-Lastly, we compare our proposed methods to two iterative algorithms optimizing the ML formulation in (13), namely the EM from [16] and a MP variant similar to [17], which we run both until convergence or $500$ iteration, whatever comes first.
+To evaluate the gain achieved by the semi-blind adaptions of the CGLM channel estimators, we also compare to the pilot-only GMM [30] and VAE [32].
+Lastly, we compare our proposed methods to two iterative algorithms optimizing the ML formulation in (13), namely the EM from [15] and a MP variant similar to [16], which we run both until convergence or $500$ iterations, whichever comes first.
 VI Numerical Simulations
@@
 (b) Measurement Data (Sec. II-B)
-Figure 3: NMSE over the SNR for given channel estimations based on $N=200$ observations including one pilot per user in a $J=8$ user scenario using (a) the spatial channel model and (b) measurement data.
+Figure 4: NMSE over the SNR for given channel estimations in a $J=8$ user and $M=64$ antennas scenario based on $N=200$ observations, including $N_{p}=J$ pilots and Gaussian symbols, using (a) the spatial channel model and (b) measurement data.
+Figure 5: NMSE over the number of antennas for given channel estimations in a $J=M/8$ user scenario based on $N=25J$ observations, including $N_{p}=J$ pilots and Gaussian symbols at SNR =0=0 dB, using the spatial channel model.
 To evaluate our proposed methods, we use channel realizations, which are normalized with $\mathbb{E}\left[\|{\bm{h}}\|^{2}\right]=M$.
-Thus, we can define the SNR $=\frac{1}{\sigma^{2}}$.
-Further, the normalized MSE (NMSE) defined as
+Thus, we can define the SNR as $\sigma^{-2}$.
+Further, the normalized MSE (MSE) defined as
 $\text{NMSE}=\frac{1}{ML}\sum_{\ell=1}^{L}\|{\bm{h}}_{\ell}-\hat{{\bm{h}}}_{\ell}\|^{2},$ (63)
-is used to characterize the performance of the estimators based on $L=10^{3}$ unseen channel samples stemming from the channel models detailed in Sections II-A and II-B.
+is used to characterize the estimators’ performances based on $L=10^{3}$ unseen channel samples stemming from the channel models detailed in Sections II-A and II-B.
+We use $\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ with $T=1.5\cdot 10^{5}$ training samples from the respective channel model to train the GMM and VAE, where we set the number of components to $K=64$ and the latent dimension to $Z=32$, respectively.
+Further, in the case of the VAE we allow non-zero values for $\bm{\mu}_{\bm{\theta}}({\bm{z}})$, as we use the circulant approximation for the spatial channel model and the block-Toeplitz property is not perfectly fulfilled for the measurement data, due to hardware imperfections.
+The VAE’s remaining implementation details are the same as provided in [32].
 The assumption of spatial uncorrelated channels, as used in (23), only holds for the spatial channel model of Section II-A.
-For the case of the measurement campaign described in Section II-B, we approximate the noise covariance matrix (24) as
+For the case of the measurement campaign described in Section II-B, we approximate the noise covariance matrix in (22) as
 ${\bm{C}}_{\tilde{{\bm{n}}}}\approx\sigma^{2}\frac{J}{M}\mathbf{I}_{M}.$ (64)
-We use =$\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ with $T=1.5\cdot 10^{5}$ training samples from the respective channel model to train the GMM and VAE, where we set the number of components to $K=64$ and the latent dimension to $Z=32$, respectively.
-Further, in the case of the VAE we allow non-zero values for $\bm{\mu}_{\bm{\theta}}({\bm{z}})$, as we use the circulant approximation for the spatial channel model and the block-Toeplitz property is not perfectly fulfilled for the measurement data, due to hardware imperfections.
+Using the $T$ training samples to empirically approximate the projected noise covariance matrix (22) did not result in a different performance compared to using the approximation (64).
 The “s-cov” variants (“sub.
@@
 s-cov”) utilize the same training samples.
-The number of BS antenna is set to $M=64$, cf.
-Sections II-A and II-B, serving $J=8=M/8$ number of users, a representative operating point [37, Chap.
+For most of the simulations, the number of BS antennas is set to $M=64$, cf.
+Sections II-A and II-B, serving $J=8=M/8$ number of users, a representative operating point [38, Chap.
 1.3.3].
-Further, the number of snapshots is set to $N=200$, if not stated otherwise, corresponding to a scenario that allows high channel dispersion and high mobility, e.g., up to $135$ kph, c.f.
-[37, Chap.
+Further, if not stated otherwise, the number of snapshots is set to $N=200$, corresponding to a scenario that allows high channel dispersion and mobility, e.g., up to $135$ kph, c.f.
+[38, Chap.
 2.1].
-The sent symbols during data transmission generally stem from a discrete constellation, e.g., QPSK, $16$-QAM.
-For this work, we utilize Gaussian symbols with $x_{j}(n)\sim\mathcal{N}_{\mathbb{C}}(0,P_{j}=1/J)$ such that $\sum_{j=1}^{J}P_{j}=1$.
-Using a continuous symbol constellation has a negligible effect on the results of the simulations, as also previously observed in [16].
+In Fig.
+3, the channel estimation performances are compared for different symbol constellations using the proposed GMM based semi-blind channel estimation variants.
+Additionally, we compare to the purely pilot-based GMM channel estimator with $N_{p}=J$ and $N_{p}=N$ number of pilots, where the latter assumes all sent symbols to be known by the receiver.
+We utilize Gaussian symbols with $x_{j}(n)\sim\mathcal{N}_{\mathbb{C}}(0,P_{j}=1/J)$ such that $\sum_{j=1}^{J}P_{j}=1$ as well as QPSK and 1616QAM.
+In the latter two cases, also a decision directed GMM is evaluated, which decodes the sent symbols based on a pilot-based GMM channel estimate and uses the decoded symbols as additional pilots in a second stage.
+For the decoding, we utilize LMMSE equalization with a subsequent mapping to the closed constellation point.
+Firstly, we see no qualitative difference between the different symbol constellations.
+Using a continuous symbol constellation has a negligible effect on the simulation results, as also previously observed in [15] and, hence, for the rest of this work, we utilize Gaussian symbols.
+Further, we see in Fig.
+3 an improvement from the pilot-only GMM to our proposed semi-blind variants, showcasing again the advantage of utilizing the payload symbols in channel estimation.
+The full pilot GMM using $N_{p}=N$ pilot symbols achieves the SNR improvement as shown in Section II.
+The decision directed GMM based estimator saturates for high SNR values, making it an inferior semi-blind strategy.
+(a) Spatial Channel Model (Sec. II-A)
+(b) Measurement Data (Sec. II-B)
+Figure 6: NMSE over the number of observations for given channel estimations in a $J=8$ user and $M=64$ antennas scenario based on NN observations, including $N_{p}=J$ pilots and Gaussian symbols at SNR =0=0 dB, using (a) the spatial channel model and (b) measurement data.
 Fig.
-3a and Fig.
-3b show the performance of the different channel estimation methods with respect to the SNR for the spatial channel model (cf.
+4a and Fig.
+4b show the different channel estimators’ performances with respect to the SNR for the spatial channel model (cf.
 Section II-A) and measurement data (cf.
@@
 One can see that the semi-blind methods utilizing the CGLMs perform the best across the whole SNR.
-The projected variant slightly outperforms its subspace counterpart for most SNR values, which follows the derivations in Section III.
+The projected variants of the CGLMs slightly outperform their subspace counterpart for most SNR values, which follows the derivations in Section III.
 Interestingly, the order of the semi-blind GMM and semi-blind VAE depends on the utilized channel model.
 In Fig.
-3a, the projected GMM and VAE show both the best overall result, whereas in Fig.
-3b, the projected GMM outperforms all other estimators.
+4a, the projected GMM and VAE show both the best overall result, whereas in Fig.
+4b, the projected GMM outperforms the VAE-based estimator.
 Additionally, we see that in Fig.
-3a, the subspace VAE outperforms the subspace GMM and in Fig.
-3b, the results are vice versa.
+4a, the subspace VAE outperforms the subspace GMM, and in Fig.
+4b, the results are vice versa.
 This ordering follows the ordering of the standalone version, where the plain GMM is better than the plain VAE in the case of the measurement data and worse for the spatial channel model.
+This comes from the fact that the Toeplitz assumption at the VAE may not be fulfilled due to impairments in the real measurement data.
 For high SNR values, all semi-blind variants approach each other except the EM and MP methods.
 In Fig.
-3a, the EM and MP drastically improve from $15$dB to $20$dB showing similar performance at $20$dB as the other semi-blind methods, whereas, in Fig.
-3b, they show inferior results also for high SNR.
+4a, the EM and MP drastically improve from 1515dB to 2020dB, showing similar performance at 2020dB as the other semi-blind methods, whereas, in Fig.
+4b, they show inferior results also for high SNR.
 The CGLM-based approaches keep a slight advantage even for high SNR values, which can be attributed to the fact that prior information is beneficial even for high SNR.
 A notable observation is that, in the mid-SNR range, the semi-blind CGLM variants outperform all related estimators by roughly $3$ dB.
-(a) Spatial Channel Model (Sec. II-A)
-(b) Measurement Data (Sec. II-B)
-Figure 4: NMSE over the number of observations for given channel estimations based on $N$ observations including one pilot per user at SNR $=0$ dB in a $J=8$ user scenario using (a) the spatial channel model and (b) measurement data.
-For our proposed strategies, the accuracy of the estimated subspace $\mathrm{range}(\hat{{\bm{V}}})$ influences the performance and, hence, the NMSE depends on the number of snapshots $N$ as shown in Fig.
-4.
-We see that for an increasing number of snapshots, the NMSE of our proposed methods decreases.
-In the case of the spatial channel model (Fig.
-4a) the projected GMM and VAE perform best for low numbers of snapshots, where the standalone VAE surpasses all other methods for $N=20$.
-Additionally, we observe in Fig.
-4a that for high $N$, the subspace VAE becomes the best of all considered methods.
 In Fig.
-4b, we observe again that for the measurement data, the semi-blind GMM variants perform the best, where for high $N$ the subspace GMM and low $N$ the projected GMM outperforms all other methods.
-Again, for less than $30$ snapshots, the semi-blind methods are outperformed by the standalone GMM and VAE due to inaccuracies in estimating the subspace with a low number of payload data symbols.
-For both utilized channel models, the subspace variant of the superior CGLM surpasses the projected variant for high numbers of snapshots converging to a lower error level.
-Thus, in practice, where, in general, uncorrelated Rayleigh fading is not the case, there are cases where the subspace CGLM outperforms its projected counterpart.
+5 we evaluate the respective estimators across different numbers of antennas at the BS.
+In order to make the comparison fair, we keep the number of users at a constant ratio of $J=M/8$ and set the number of snapshots to $N=25J$.
+With this, we only see an improvement for higher numbers of antennas for the CGLMs, while the other methods roughly keep their performance.
+Due to the constant parameter ratio, the VAE variants’ drastic performance gains for a higher number of antennas solely come from the VAE’s strong capability of modeling high-dimensional data.
+[…]
 A critical decrease in performance can be observed for the EM and MP algorithms.
@@
 This is because both methods optimize the joint ML formulation in (13), where the optimization of the second term becomes dominant for a high number of snapshots.
-Hence, the impact of the pilot observations relevant to estimating the phase of the channel vanishes.
+Hence, the pilot observation’s impact, which is relevant to estimating the channel phase, vanishes.
 (a) Spatial Channel Model (Sec. II-A)
 (b) Measurement Data (Sec. II-B)
-Figure 5: NMSE over the number of users for given channel estimations based on $N=200$ observations including one pilot per user at SNR $=0$ dB using (a) the spatial channel model and (b) measurement data.
-The dimension of the subspace $\mathrm{range}(\hat{{\bm{V}}})$ directly influences the estimation quality of the proposed methods as shown in Fig.
-5a and Fig.
-5b.
-For example, in the extreme case where the number of users in the system is equal to the number of BS antennas ($J=M$), the solution to (56) becomes $\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}=\mathbf{I}$ and, hence, as the number of users in the system increases all semi-blind estimators approach their respective purely pilot based version.
-We restrict our simulations within the interval of $J\leq M/4=16$, which is said to be the preferred operating regime in massive MIMO [37, Chap.
-1.3.3], and set the number of snapshots to $N=200$.
-In the case of a single user, all semi-blind variants exhibit similar performance, except for the subspace sample covariance estimator, the subspace VAE, and in the case of the spatial channel model (Fig.
-5a) the subspace GMM.
-For all other considered numbers of users, the proposed projected CGLM methods outperform all other channel estimators.
-Additionally, for the spatial channel model in Fig.
-5a, the subspace GMM also shows inferior results to the other CGLM-based methods for all numbers of users.
+Figure 7: NMSE over the number of users for given channel estimations in a $M=64$ antennas scenario based on $N=200$ observations, including $N_{p}=J$ pilots and Gaussian symbols at SNR =0=0 dB, using (a) the spatial channel model and (b) measurement data.
+[…]
 Overall, we can conclude that the proposed semi-blind CGLMs show superior channel estimation performance across all different setups.
 Depending on the used channel model, either the semi-blind GMMs or the semi-blind VAEs result in slightly better NMSE, where only in the case of the spatial channel model, the subspace GMM shows slightly worse performance compared to the other proposed methods.
-Moreover, the projected CGLMs outperform their respective subspace counterpart for most simulated operating points, showing the superiority of the proposed projection method.
+Moreover, the projected CGLMs outperform their respective subspace counterpart for most simulated operating points, showing the proposed projection method’s superiority.
 VII Conclusion
 This work presented a novel semi-blind channel estimation technique based on the class of CGLMs.
-To this end, two methods are discussed that incorporate subspace knowledge about the channel into the well-known LMMSE estimator.
+To this end, two methods that incorporate subspace knowledge into the well-known LMMSE estimator are discussed.
 Both methods exploit the estimated subspace derived from the dominant eigenvectors of sample covariance matrices constructed using the received symbols.
-A theoretical analysis of the methods showed the superior estimation quality of the proposed projection-based estimator for uncorrelated Rayleigh fading channels.
+A theoretical analysis of the methods showed the proposed projection-based estimator’s superior estimation quality for uncorrelated Rayleigh fading channels.
 Secondly, we showed how two examples from the class of CGLMs, i.e., the GMM and VAE, can be used to parameterize these estimators.
-Extensive simulations based on real-world measurement and spatial channel model data demonstrated the superior estimation performance of the proposed methods compared to standard semi-blind channel estimators.
+Extensive simulations based on real-world measurement and spatial channel model data demonstrated the proposed methods’ superior estimation performances compared to standard semi-blind channel estimators.
 -A MSE of Projected LMMSE
@@
 $\mathrm{MSE}$ $=\mathbb{E}\left[\|{\bm{h}}-\hat{{\bm{h}}}\|^{2}\right]=\mathbb{E}\left[\mathrm{tr}\left(\left({\bm{h}}-\hat{{\bm{h}}}\right)\left({\bm{h}}^{\mathrm{H}}-\hat{{\bm{h}}}^{\mathrm{H}}\right)\right)\right]$ (65)
-$=\mathbb{E}\left[\mathrm{tr}({\bm{h}}{\bm{h}}^{\mathrm{H}})-2\mathrm{tr}({\bm{h}}{\bm{y}}^{\mathrm{H}}{\bm{W}})+\mathrm{tr}({\bm{W}}{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}})\right].$ (64)
+$=\mathbb{E}\left[\mathrm{tr}({\bm{h}}{\bm{h}}^{\mathrm{H}})-2\mathrm{tr}({\bm{h}}{\bm{y}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}})+\mathrm{tr}({\bm{W}}{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}})\right].$ (66)
 For the case of the projected LMMSE the second term in (66) can be rewritten as
@@
 $\quad\quad=\mathbb{E}\bigg[\mathrm{tr}\bigg({\bm{C}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}+\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right)$
-×$\quad\quad\quad\times\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\bigg{)}\bigg{]}$ (71)
+$\quad\quad\quad\times\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\bigg)\bigg]$ (73)
 $\quad\quad=\mathrm{tr}\left({\bm{C}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\right),$ (74)
-where we assume that $\mathbb{E}\left[\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right]=\sigma^{2}\frac{J}{M}\mathbf{I}_{M}$. From this the overall MSE in (28) follows directly.
+where we assume that $\mathbb{E}\left[\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right]=\sigma^{2}\frac{J}{M}\mathbf{I}_{M}$. From this, the overall MSE in (26) follows directly.
 -B MSE of Subspace LMMSE for Rayleigh Fading
-In the case of uncorrelated Rayleigh fading the subspace LMMSE filter is given as
+In the case of uncorrelated Rayleigh fading, the subspace LMMSE filter is given as
 ${\bm{W}}_{\mathrm{sub}}=\frac{1}{1+\sigma^{2}}{\bm{V}}{\bm{V}}^{\mathrm{H}}.$ (75)
@@
 $\quad\quad=\frac{1}{(1+\sigma^{2})^{2}}(M+J\sigma^{2}).$ (85)
-The overall MSE of the subspace variant for ${\bm{C}}=\mathbf{I}_{M}$ is then
+The overall subspace variant’s MSE for ${\bm{C}}=\mathbf{I}_{M}$ is then
 $\mathrm{MSE}_{\mathrm{iid}}^{\mathrm{sub}}$ $=M-2\frac{1}{1+\sigma^{2}}M+\frac{1}{(1+\sigma^{2})^{2}}(M+J\sigma^{2})$ (86)
 $=\frac{\sigma^{2}(M\sigma^{2}+J)}{(1+\sigma^{2})^{2}}.$ (87)
+-C Derivation of Bayesian Cramer-Rao Bounds
+The BCRB, which is the lower bound on the MSE of any estimator $\hat{{\bm{h}}}({\bm{Y}})$, is given as [48]
+$\mathbb{E}\left[\|\hat{{\bm{h}}}({\bm{Y}})-{\bm{h}}\|^{2}\right]\geq\mathrm{tr}\left(\left[\left(\mathbb{E}_{{{\bm{h}}}}[\mathcal{I}({{{\bm{h}}}})]+\bm{J}_{p}\right)^{-1}\right]_{:M,:M}\right),$ (88)
+where $\mathcal{I}({{{\bm{h}}}})$ is the Fisher information matrix corresponding to the inverse of the unbiased CRB, and
+$\bm{J}_{p}$ $=\mathbb{E}_{\bm{h}}\left[\frac{\partial\log p({\bm{h}})}{\partial{\bm{h}}^{\mathrm{H}}}\frac{\partial\log p({\bm{h}})}{\partial{\bm{h}}}\right]$ (89)
+$=\mathrm{blckdiag}({{\bm{C}}}_{1},\dots,{{\bm{C}}}_{J})^{-1},$ (90)
+We denote with $\cdot]_{:M,:M}$ in (88) the part of the overall BCRB matrix corresponding to the upper left block of size $M\times M$, where we assume without loss of generality to index the users such that the user of interest is the first one.
+For the case of ${\bm{C}}_{j}=\mathbf{I}_{M}$ for all $j\in\{1,\dots,J\}$ the prior information $\bm{J}_{p}$ is given as
+$\bm{J}_{p}=\mathbf{I}_{JM}.$ (91)
+For the deterministic BCRB, we use the formulation derived in [40], which results in
+$\mathbf{BCRB}_{\mathrm{iid,d}}$
+$\quad=\sigma^{2}\left(({\bm{X}}^{*}{\bm{X}}^{\mathrm{T}})\otimes{\bm{P}}_{\bm{H}}^{\perp}+({\bm{P}}^{*}{\bm{P}}^{\mathrm{T}})\otimes{\bm{P}}_{\bm{H}}+\sigma^{2}\mathbf{I}_{JM}\right)^{-1},$ (92)
+with ${\bm{P}}_{\bm{H}}^{\perp}=\mathbf{I}_{M}-{\bm{P}}_{\bm{H}}$.
+Now using the Fisher information matrix for the stochastic CRB as provided in [15], we introduce $\bar{{\bm{h}}}=[\Re\{{\bm{h}}_{1}^{\mathrm{T}}\},\Im\{{\bm{h}}_{1}^{\mathrm{T}}\},\dots,\Re\{{\bm{h}}_{J}^{\mathrm{T}}\},\Im\{{\bm{h}}_{J}^{\mathrm{T}}\}]^{\mathrm{T}}$, which gives us the prior information as
+$\bm{J}_{p}$ $=\mathbb{E}_{\bar{{\bm{h}}}}\left[\frac{\partial\log p(\bar{{\bm{h}}})}{\partial\bar{{\bm{h}}}^{\mathrm{T}}}\frac{\partial\log p(\bar{{\bm{h}}})}{\partial\bar{{\bm{h}}}}\right]$ (93)
+$=2\,\mathrm{blckdiag}(\bar{{\bm{C}}}_{1},\dots,\bar{{\bm{C}}}_{J})^{-1},$ (94)
+with
+$\bar{{\bm{C}}}_{j}=\begin{bmatrix}\Re\{{\bm{C}}_{j}\}&-\Im\{{\bm{C}}_{j}\}\\ \Im\{{\bm{C}}_{j}\}&\Re\{{\bm{C}}_{j}\}\end{bmatrix}$. (95)
+Thus, the BCRB for uncorrelated Rayleigh fading is given as
+$\mathbf{BCRB}_{\mathrm{iid,s}}=\left(\left(\frac{2N_{p}}{J\sigma^{2}}+2\right)\mathbf{I}_{2JM}+{\bm{R}}\right)^{-1},$ (96)
+where $\bm{R}$} denotes the part depending on the payload symbols defined in [15].
+-D Proof of Theorem 2
+To this end, let us consider the MSE for a fixed channel matrix $\bm{H}$}, where the expectation is only taken over the noise and Gaussian symbol realizations as
+$\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]$
+$\quad=\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|{\bm{W}}_{\mathrm{proj}}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}){\bm{y}}\|^{2}]$ (97)
+$\quad\leq\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})\mathbb{E}_{{\bm{X}}}\left[(\mathrm{tr}({\bm{A}})+\mathbb{E}_{{\bm{N}}}[\mathrm{tr}({\bm{B}})])\right]$ (98)
+with
+$\mathrm{tr}({\bm{A}})$ $=\mathrm{tr}((\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}){\bm{h}}{\bm{h}}^{\mathrm{H}}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}))$ (99)
+$=\mathrm{tr}((\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}){\bm{h}}{\bm{h}}^{\mathrm{H}}(\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}))$ (100)
+$=\|(\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}){\bm{h}}\|^{2}$ (101)
+$=\|(\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}){\bm{V}}{\bm{s}}\|^{2}$ (102)
+$=\sum_{i\leq J}\sum_{j>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}s_{i}|^{2}$ (103)
+$\leq|s_{\max}|^{2}\sum_{j\leq J}\sum_{i>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}$ (104)
+$=|s_{\max}|^{2}\sum_{i\leq J}\sum_{j>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}$ (105)
+where ${\bm{h}}={\bm{V}}{\bm{s}}$ with ${\bm{s}}=[s_{1},\dots,s_{J}]^{\mathrm{T}}$ is the channel of interest’s decomposition in terms of $\bm{V}$}, and $s_{\max}=\max_{i}s_{i}$. Furthermore, we have
+$\mathbb{E}_{\bm{N}}[\mathrm{tr}({\bm{B}})]$ $=\mathbb{E}_{\bm{N}}[\mathrm{tr}((\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}){\bm{n}}{\bm{n}}^{\mathrm{H}}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}))]$ (106)
+$=\sigma^{2}(\mathrm{tr}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}+{\bm{V}}{\bm{V}}^{\mathrm{H}})-2\mathrm{tr}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}{\bm{V}}{\bm{V}}^{\mathrm{H}}))$ (107)
+$=2\sigma^{2}\left[J-\sum_{i,j\leq J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}\right]$ (108)
+$=2\sigma^{2}\sum_{j\leq J}\sum_{i>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}$ (109)
+Thus,
+$\mathbb{E}_{{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]$
+$\quad\leq\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})\left(|s_{\max}|^{2}+2\sigma^{2}\right)\sum_{j\leq J}\sum_{i>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}.$ (110)
+Now, let us consider the following corollary from [49, Corollary 4.1] about the inner product of eigenvectors of sample and true covariance matrices.
+Corollary 2.1.
+Let us denote with $\lambda^{\prime}_{j}$ and ${\bm{v}}_{j}$ the $i$-th eigenvalue and eigenvector of the true covariance matrix, respectively.
+Then, for any weights $w_{ji}$, which are non-zero when $\lambda^{\prime}_{i}\neq\lambda^{\prime}_{j}$ and $\mathrm{sgn}(\lambda^{\prime}_{i}-\lambda^{\prime}_{j})2\lambda^{\prime}_{i}>\mathrm{sgn}(\lambda^{\prime}_{i}-\lambda^{\prime}_{j})(\lambda^{\prime}_{i}+\lambda^{\prime}_{j})$, and real $\varepsilon>0$
+$P\left(\sum_{j\leq J}\sum_{i>J}w_{ji}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}\leq\varepsilon\right)$
+$\quad\geq 1-\sum_{j\leq J}\sum_{i>J}\frac{4k^{2}_{i}w_{ji}}{N\varepsilon(\lambda^{\prime}_{j}-\lambda^{\prime}_{i})^{2}},$ (111)
+where $k_{i}^{2}=\mathbb{E}[\|{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{v}}_{i}\|^{2}]-\lambda_{i}^{\prime,2}$.
+In our given problem, the true covariance matrix is given as ${\bm{C}}_{{\bm{y}}\mid{\bm{H}}}={\bm{H}}{\bm{H}}^{\mathrm{H}}/J+\sigma^{2}\mathbf{I}_{M}$.
+It is then trivial to see that $\lambda^{\prime}_{i}=\sigma^{2},\forall i>J$.
+Furthermore it is shown that [49, Corollary 4.3]
+$k_{i}^{2}=\lambda^{\prime}_{i}\left(\lambda^{\prime}_{i}+\mathrm{tr}(C_{{\bm{y}}\mid{\bm{H}}})\right).$ (112)
+Hence, we can conclude
+$P\left(\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]\leq\varepsilon\right)$
+$\geq 1-\frac{4k^{2}(M-J)\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})}{J^{2}N\varepsilon}\sum_{j\leq J}\frac{|s_{\max}|^{2}+2\sigma^{2}}{\lambda_{j}^{2}},$ (113)
+with $\lambda_{j}=(\lambda^{\prime}_{j}-\sigma^{2})/J$ for all $j\in\{1,\dots,J\}$ being the eigenvalues of ${\bm{H}}{\bm{H}}^{\mathrm{H}}$ and
+$k^{2}=\sigma^{2}\left(\sigma^{2}+\mathrm{tr}(C_{{\bm{y}}\mid{\bm{H}}})\right).$ (114)